### v1

# GUI-World: A Dataset for GUI-oriented Multimodal LLM-based Agents

Abstract
Recently, Multimodal Large Language Models (MLLMs) have been used as agents to control keyboard and mouse inputs by directly perceiving the Graphical User Interface (GUI) and generating corresponding code. However, current agents primarily exhibit excellent understanding capabilities in static environments and are predominantly applied in relatively simple domains, such as Web or mobile interfaces. We argue that a robust GUI agent should be capable of perceiving temporal information on the GUI, including dynamic Web content and multi-step tasks. Additionally, it should possess a comprehensive understanding of various GUI scenarios, including desktop software and multi-window interactions. To this end, this paper introduces a new dataset, termed GUI-World, which features meticulously crafted Human-MLLM annotations, extensively covering six GUI scenarios and eight types of GUI-oriented questions in three formats. We evaluate the capabilities of current state-of-the-art MLLMs, including ImageLLMs and VideoLLMs, in understanding various types of GUI content, especially dynamic and sequential content. Our findings reveal that ImageLLMs struggle with dynamic GUI content without manually annotated keyframes or operation history. On the other hand, VideoLLMs fall short in all GUI-oriented tasks given the sparse GUI video dataset. Based on GUI-World, we take the initial step of leveraging a fine-tuned VideoLLM as a GUI agent, demonstrating an improved understanding of various GUI tasks. However, due to the limitations in the performance of base LLMs, we conclude that using VideoLLMs as GUI agents remains a significant challenge. We believe our work provides valuable insights for future research in dynamic GUI content understanding. The code and dataset are publicly available at our project homepage: https://gui-world.github.io/.
1 Introduction
Multimodal Large Language Models (MLLMs), such as GPT-4V(ision) [1] and LLaVA [2], have significantly contributed to the development of the visual-text domain [3]. These models bring forth innovative solutions and paradigms for traditional visual tasks, including visual reasoning [4], medical image interpretation [5, 6], and applications in embodied agents [7]. One particularly promising area is Graphical User Interface (GUI) understanding, which holds significant potential for real-world applications, such as webpage comprehension [8, 9] and navigation by GUI agents [10, 11, 12]. The key challenges of GUI understanding are twofold: effective GUI agents are expected to (1) possess a deep understanding of GUI elements, including webpage icons, text identified through Optical Character Recognition (OCR), and page layouts, and (2) exhibit an exceptional ability to follow instructions within GUI contexts, such as conducting searches through search engines.
Despite significant progress, as illustrated in Table 1, existing works suffer from the following limitations: (1) Most studies predominantly focus on the static features of GUI scenarios, neglecting the need for MLLMs to effectively process sequential information and dynamic operations. For instance, an agent’s task performance can be disrupted by unexpected elements such as pop-up advertisements, underscoring a gap in handling dynamic sequential tasks. (2) Current research is typically restricted to Web-based environments, which limits the models’ generalization and robustness. For instance, GUI agents may need to operate across diverse platforms such as Windows, macOS, Linux, iOS, Android, and XR environments. Additionally, operations may sometimes involve multiple windows. Therefore, expanding the scope of research to encompass these varied environments will enhance the adaptability and effectiveness of GUI agents.
To mitigate these gaps, this paper introduces GUI-World, a comprehensive dataset containing over 12,000 GUI videos, specifically designed to evaluate and enhance the capabilities of GUI agents. This dataset encompasses a wide range of GUI scenarios, including popular websites, desktop and mobile applications across various operating systems, multi-window interactions, as well as XR environments. The data collection process involves sourcing GUI videos from screen recordings and instructional videos on YouTube. Subsequently, we utilize an Human-MLLM collaborative approach to generate a diverse set of questions and instructions and finally construct GUI-World.
| Dataset | Size | Sem. | VL | Video | Env Type | Task Coverage | Task | |||||
| Web. | Mob. | Desk. | XR | Seq. | Cro. | Dyn. | ||||||
| Rico [13] | 72,219 | Low | ✔ | ✔ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | UI Code/Layout Generation |
| MetaGUI [14] | 1,125 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Mobile Navigation |
| UGIF [15] | 523 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | UI Grounded Instruction Following |
| AITW [16] | 715,142 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| Ferret-UI [17] | 123,702 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Grounding & Understanding |
| MiniWoB++ [18] | 100 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | ✘ | Web Navigation |
| WebArena [19] | 812 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| Mind2Web [20] | 2,350 | Both | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| OmniAct [21] | 9,802 | Low | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | Code Generation |
| MMINA [22] | 1,050 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✔ | ✘ | Web Navigation |
| AgentStudio [23] | 304 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| OSWorld [24] | 369 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| GUI-World (Ours) | 12,379 | Both | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | GUI Understanding |
| Instruction Following | ||||||||||||
Likewise, we also establish a comprehensive benchmark for GUI understanding, which encompasses seven mainstream MLLMs, three keyframe selection strategies, six GUI scenarios, and a diverse array of queries in multiple-choice, free-form, and conversational formats, aiming to provide a thorough evaluation of the MLLMs’ GUI-oriented capabilities. As shown in Figure 2, the assessment results indicate that most MLLMs struggle with GUI-World, highlighting their limited dynamic understanding of graphical interfaces and underscoring the need for further enhancement.
Leveraging this dataset, we take the first step of fine-tuning a Video GUI Agent proficient in dynamic and sequential GUI tasks, which results in significant improvements in the general capabilities of GUI agents, thereby demonstrating the utility and effectiveness of GUI-World. Additionally, we delve into discussing various factors critical to GUI understanding, including the integration of textual information, the number of keyframes, and image resolutions.
Overall, the key contributions of this paper are three-fold:
A New Dataset. We propose GUI-World, a comprehensive GUI dataset comprising over 12,000 videos specifically designed to assess and improve the GUI understanding capabilities of MLLMs, spanning a range of categories and scenarios, including desktop, mobile, and extended reality (XR), and representing the first GUI-oriented instruction-tuning dataset in the video domain.
A Novel Model. Based on GUI-World, we propose GUI-Vid, a GUI-oriented VideoLLM with enhanced capabilities to handle various and complex GUI tasks. GUI-Vid shows a significant improvement on the benchmark and achieves results comparable to the top-performing models.
Comprehensive Experiments and Valuable Insights. Our experiments indicate that most existing MLLMs continue to face challenges with GUI-oriented tasks, particularly in sequential and dynamic GUI content. Empirical findings suggest that improvements in vision perception, along with an increase in the number of keyframes and higher resolution, can boost performance in GUI-oriented tasks, thereby paving the way for the future of GUI agents.
2 GUI-World: A Comprehensive Dataset for GUI Understanding
2.1 Overview
We introduce GUI-World, a comprehensive dataset covering six GUI scenarios including video, human-annotated keyframes, as well as detailed captions and diverse types of QA produced by our data curation framework, aiming at benchmarking and enhancing the general GUI-oriented capabilities. These GUI scenarios encompass desktop operating systems (e.g., macOS, Windows) and mobile platforms (e.g., Android and iOS), websites, software, and even extended-range technologies (XR) (e.g., GUI in Apple Vision Pro [25]). Discussion for each scenario is in subsection A.1.
As illustrated in Figure 3, the development of GUI-World is structured around a two-stage process. Details regarding video and query statistics are provided in Table 2, which includes distributions of the number of keyframes, video lengths, and the lengths of queries and their corresponding golden answers, as displayed in Figure 4. Refer to Figure 5 and Appendix F for case study.
2.2 GUI Video Collection and Keyframe Annotation Process
We describe the pipeline for collecting screen recordings from student workers and GUI-related instructional videos from YouTube for GUI-World and the procedures followed to convert these videos into keyframe sequences.
| Category | Total Videos | Free-form | MCQA | Conversation | Total Frame. (Avg.) | Avg. Anno. |
| Software | 4,720 | 27,840 | 9,440 | 9,440 | 23,520 (4.983) | 7.558 |
| Website | 2,499 | 14,994 | 4,998 | 4,998 | 15,371 (6.151) | 6.862 |
| IOS | 492 | 2,952 | 984 | 984 | 2,194 (4.459) | 7.067 |
| Multi | 475 | 2,850 | 950 | 950 | 2,507 (5.277) | 7.197 |
| XR | 393 | 2,358 | 786 | 786 | 1,584 (4.030) | 10.970 |
| Android | 3,800 | 15,199 | 7,600 | 7,600 | 38,000 (10.000) | - |
| Summary | 12,379 | 76,673 | 24,758 | 24,758 | 83,176 (6.719) | 7.463 |
A significant portion of our video data is derived from screen recordings executed by student workers, which can directly reflect real-life GUI usage scenarios. A typical video collection scenario involves assigning a student worker a specific software task. The student begins by familiarizing themselves with the software, followed by recording a series of operations in a short video clip, such as “Sign up”, “Sign in”, “Create a New Page”, and “Invite Other Collaborators” in the software “Notion111https://www.notion.so/”.
Despite the high fidelity of these manually recorded videos, we encounter several challenges: (1) Student workers often require substantial time to acquaint themselves with professional software (e.g., MATLAB, Adobe After Effects (Ae)), which can hinder the progress of data collection. (2) The videos may lack comprehensiveness, typically capturing only commonly used operations and overlooking rarer functions crucial for dataset completeness. To address these issues, we also source videos from social media platforms that host a diverse array of GUI-related content. Specifically, we download tutorial videos from YouTube—given its prevalence as a video-sharing platform—because they richly detail various GUI operations. These videos are then segmented into shorter clips, each representing a distinct sequence of operations.
The subsequent step involves annotating these video clips with keyframes and textual descriptions of each keyframe using custom-designed annotation software. Although several algorithms exist for keyframe extraction [26, 27, 28, 29], they typically underperform with GUI videos where changes between frames might be minimal (e.g., a slight movement in the mouse cursor). To ensure high-quality datasets, we therefore perform manual extraction of these keyframes. Each keyframe is meticulously annotated to include details such as the operation performed, the purpose between two keyframes, the software or website used, mouse actions (e.g., scroll, click), and keyboard inputs (e.g., copy (Ctrl + C), paste (Ctrl + V), specific input). We detail our annotation process in subsection A.3.
2.3 GUI Tasks Generation from Human-MLLM Collaboration
Drawing insights from prior research [30, 31, 32, 33, 34], we develop a Human-MLLM collaboration pipeline to annotate captions and diverse types of QA specifically tailored for GUI comprehension. The process involves inputting an instructional prompt, a comprehensive description, key information (e.g., system or application), and a sequence of human-annotated keyframes into GPT-4V. As depicted in Table 3, GUI-World features an array of question types, as detailed in follows:
Detailed and Summarized Captioning: This task challenges basic GUI knowledge and multimodal perception, also addressing the deficiency of detailed GUI content in video-caption pairs. Initially, GPT-4V generates two distinct descriptions for each video: one concentrating on fine-grained details and the other on the overall image sequences. Furthermore, GPT-4V provides a succinct summary, highlighting core operations and overarching objectives in the video.
Static GUI Content: This task challenges MLLM with textual, layout, and iconographic analysis of static GUI content. We instruct GPT-4V to generate free-form queries with a golden answer concerning static GUI elements or specific scenes that recur in more than two keyframes, ensuring their consistent presence in the video. Additionally, GPT-4V also crafts QA pairs that evaluate inferential skills in static content, focusing on interrelations among icons or textual information.
Dynamic and Sequential GUI Content: This task concentrates on temporal content in GUI video, such as dynamically changing interfaces, and aims to elucidate the sequential information and reasoning chains within GUI content. We direct GPT-4V to identify consistently changing elements to create queries for dynamic content. Moreover, predictive tasks are formulated on order and temporal relation in provided sequential images, challenging agents to anticipate future events or states.
| T. | Question | Examples |
| Caption | Detailed | Q: Please provide a detailed description of what occurs throughout these sequential GUI images. |
| Description | A: The video shows a user taking the 16 Personalities test on a Windows desktop using the Edge browser… | |
| Summarized | Q: Write a clear description of the video, make sure the key features are well covered. | |
| Caption | A: Creating a new IT team in Todoist by selecting industry, job function, role, team size, and inviting members. | |
| Static | Layout, | Q: What related searches are suggested on the right side of the Bing results for ’emnlp 2024’? |
| Icon Retrieval | A: The suggested related searches shown include ’emnlp 2024 miami’, ’eacl 2024 call for papers’… | |
| Textual | Q: What is the estimated time to complete the content for Week 2 of the course? | |
| Retrieval | A: The estimated time to complete the content for Week 2 of the course is 1 hour… | |
| Interrelations | Q: What is the name of the browser and the tab where the user performs the product search? | |
| in GUI Content | A: The browser is Microsoft Edge, and the user performs the product search in the eBay tab. | |
| Dynamic | Content | Q: What specific action does the user take after turning their head to the left to view the left side of the page? |
| Retrieval | A: After turning their head to the left to view the left side of the page, the user performs… | |
| Prediction | Q: Given the mouse is over ’Add NeurIPS 2024 DB Track Submission,’ what’s the likely next step? | |
| A: It would be to click on the ’Add NeurIPS 2024 Datasets and Benchmarks Track Submission’ button… | ||
| Sequential | Q: Scrolls down from the ’Moon Gravity’, which of the following cheats? A. Change Weather B. Skyfall … | |
| Reasoning | A: [[B]] |
In the last stage, human annotators will follow the guideline in subsection A.3 and carefully review the entire video and MLLM-generated QA pairs to correct inaccuracies and hallucinations, as well as supplement information for both questions and answers to make these tasks more challenging.
3 Progressive Enhancement on GUI Perception Ability
We introduce our strategy to enhance the GUI-oriented capabilities of current MLLMs on both static and dynamic GUI content. Inspired by previous studies [9, 35], we structure our methodology into two distinct fine-tuning stages, as illustrated in Figure 6. Initially, we fine-tune the MLLM on simpler tasks, such as description queries and captioning exercises, to instill a basic understanding of GUI elements. Subsequently, building on this foundation, the second stage aims to augment the MLLM’s proficiency with more complex and challenging tasks. Our fine-tuning is all based on the Supervised Fine-Tuning (SFT): , where is the input, is LLMs’ output, and denotes the model parameters that need to be optimized.
Stage-1: Learning Preliminary for GUI Content.
The initial phase focuses on aligning GUI content with a pre-trained vision encoder and a base LLM, utilizing GUI videos accompanied by detailed descriptions and captions. This phase aims to embed a robust understanding of fundamental GUI concepts and terminology within the MLLM. By engaging the model in basically captioning various GUI components, the model learns to recognize and articulate the functionalities and visual characteristics of these elements, thereby laying a solid groundwork for GUI knowledge.
Stage-2: Mastering Advanced GUI Capability.
Building on the foundational knowledge established in Stage 1, the second stage focuses on advancing the MLLM’s proficiency in interacting with GUI elements through more complex tasks. These tasks are designed to simulate real-world scenarios that the MLLM might encounter in GUI environments, which include predicting based on image sequences, engaging in conversations, retrieving both static and dynamic GUI elements, and performing reasoning tasks.
As illustrated in Figure 6, We employ the two-stage training architecture utilizing VideoChat2 [35] as our foundational model. Initially, videos and images are encoded using the UMT-L visual encoder [36]. Subsequently, a QFormer compresses visual tokens into a smaller set of query tokens. Drawing inspiration from [37], we enhance the QFormer [38] by integrating instructions to enable it to extract visual representations pertinent to the given instructions. Additionally, we apply low-rank adaptation (LoRA [39]) to base LLM. This model is concurrently fine-tuned with the visual encoder and QFormer using a Vision-grounded Text Generation (VTG) loss: , where represents the visual tokens derived from the QFormer, and represents the text output grounded in the visual context.
4 Experiments and Analysis
4.1 Experimental Setups
Models.
We conduct evaluations on four of the most robust image-based MLLMs: GPT-4V(ision) [1], GPT-4o [40], Qwen-VL-Max [41], and Gemini-Pro-1.5 [42]. We benchmark on three keyframe selection settings: (1) Random, where frames are sampled at fixed time intervals within a video; (2) Extracted, with keyframes extracted using Katna222https://github.com/keplerlab/katna; and (3) Human, where keyframes are selected by humans during the annotation process. For the Random and Extracted settings, we input 10 frames into each MLLM, while the Human setting uses an average of 6.719 frames, as detailed in Table 2. Each model’s responses employ a three-step Chain-of-Thought (CoT) [43] process, i.e., “Describe-Analyze-Answer”, to evaluate their peak performance. Additionally, we assessed three advanced VideoLLMs—ChatUnivi [44], Minigpt4-video [45], and Videochat2 [46]—for their performance on GUI content. For detailed experimental setups are referred to Appendix C.
Evaluation Metrics.
To assess free-form questions and multiple-round conversations, we utilize the LLM-as-a-Judge methodology, which assigns a similarity score ranging from 1 to 5 between MLLM’s response and a predefined golden answer, already validated by previous studies[47, 48, 49]. For a comprehensive evaluation, we also provide BLEU [50] and BERTScore [51] in Appendix D. For multiple-choice questions, we measure performance using accuracy as the primary evaluation metric.
Textual Information Integration.
To investigate the effectiveness of integrating image-caption models to enlarge the context window for LLMs—typically employed in natural videos—and the helpfulness of GUI history content in accomplishing GUI-oriented tasks, we implement three experimental settings: Detailed Caption, Concise Caption, and Vision + Detailed Caption. GPT-4V is utilized to provide captions of these keyframes, integrating human annotators’ operational intents to more accurately describe each frame, being validated in subsection A.3.
| Models | Setting | Software | Website | XR | Multi | IOS | Android | Avg. | ||||||||
| MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |||
| ImageLLMs | Gemini-Pro-1.5 | R. | 81.7% | 3.339 | 82.6% | 3.452 | 81.2% | 3.154 | 81.2% | 2.959 | 82.0% | 3.213 | 81.6% | 3.220 | 81.7% | 3.223 |
| E. | 78.5% | 3.152 | 77.8% | 3.215 | 80.8% | 3.006 | 71.8% | 2.777 | 79.3% | 3.007 | 78.5% | 3.168 | 77.8% | 3.054 | ||
| Qwen-VL-Max | R. | 74.9% | 2.676 | 76.9% | 2.656 | 74.2% | 2.469 | 68.8% | 2.432 | 75.4% | 2.779 | 73.7% | 2.309 | 74.0% | 2.553 | |
| E. | 74.3% | 2.624 | 75.8% | 2.627 | 69.0% | 2.499 | 64.8% | 2.362 | 77.4% | 2.659 | 65.8% | 2.277 | 71.2% | 2.508 | ||
| H. | 75.8% | 2.651 | 75.5% | 2.698 | 77.6% | 2.373 | 66.9% | 2.490 | 74.3% | 2.633 | - | - | 74.0% | 2.569 | ||
| GPT-4V | R. | 81.5% | 3.589 | 80.9% | 3.648 | 80.6% | 3.200 | 75.0% | 3.452 | 82.5% | 3.614 | 78.3% | 3.515 | 79.8% | 3.503 | |
| E. | 85.1% | 3.407 | 80.1% | 3.433 | 81.8% | 2.892 | 81.9% | 3.219 | 86.4% | 3.427 | 79.9% | 3.176 | 82.6% | 3.259 | ||
| H. | 86.0% | 3.520 | 79.8% | 3.655 | 83.4% | 3.265 | 76.9% | 3.449 | 79.9% | 3.453 | - | - | 81.2% | 3.469 | ||
| D.C. | 85.0% | 3.350 | 83.1% | 3.380 | 82.3% | 3.056 | 84.2% | 3.358 | 81.6% | 2.751 | 81.7% | 3.427 | 83.0% | 3.316 | ||
| C.C | 80.7% | 3.028 | 72.2% | 3.025 | 82.8% | 2.809 | 81.3% | 3.160 | 76.5% | 2.868 | 76.4% | 2.939 | 78.3% | 2.971 | ||
| H.+D.C. | 82.5% | 3.494 | 83.2% | 3.682 | 85.9% | 3.191 | 83.9% | 3.617 | 80.9% | 3.516 | 84.9% | 3.758 | 83.5% | 3.543 | ||
| GPT-4o | H. | 86.5% | 3.644 | 83.3% | 3.740 | 84.3% | 3.285 | 81.1% | 3.654 | 83.3% | 3.558 | 90.0% | 3.561 | 84.8% | 3.573 | |
| VideoLLMs | ChatUnivi | - | 28.4% | 2.389 | 22.2% | 2.349 | 20.6% | 2.161 | 17.5% | 2.275 | 22.6% | 2.337 | 23.0% | 2.390 | 22.4% | 2.317 |
| Minigpt4Video | - | 18.9% | 1.475 | 15.3% | 1.520 | 16.3% | 1.362 | 15.4% | 1.457 | 20.1% | 1.501 | 14.6% | 1.342 | 16.8% | 1.443 | |
| VideoChat2 | - | 45.5% | 2.144 | 42.6% | 2.221 | 44.0% | 2.005 | 40.4% | 2.222 | 40.2% | 2.169 | 44.7% | 2.119 | 42.9% | 2.147 | |
| GUI-Vid | - | 59.9% | 2.847 | 54.1% | 2.957 | 55.6% | 2.764 | 52.9% | 2.861 | 51.8% | 2.773 | 53.4% | 2.572 | 54.6% | 2.796 | |
Keyframes and Resolution.
To explore the upper bound of GUI-oriented capabilities, particularly in dynamic and sequential tasks, we conduct ablation studies focusing on the impact of the number of keyframes and image resolutions. We vary the number of keyframes (8, 16) fed into GUI-Vid. Additionally, we test the effect of different image resolutions on GPT-4o, using both low and high settings, to further assess how resolution influences performance.
4.2 Empirical Results
Commercial ImageLLMs outperform Open-source VideoLLMs in Zero-shot Settings.
Commercial ImageLLMs, notably GPT-4V and GPT-4o, consistently outperform open-source VideoLLMs in zero-shot settings. As detailed in Table 4, GPT-4o exhibits superior performance across all GUI scenarios in complex tasks, reflected in its high scores in both multiple-choice and free-form queries, with an average of 84.8% and 3.573. Similarly, Gemini demonstrates strong capabilities in captioning and descriptive tasks within software and iOS environments, scoring 2.836 and 2.936, respectively, as shown in Table 13. Further analysis (Figure 7) reveals that GPT-4V excels in applications with minimal textual content and simple layouts, such as TikTok, health apps, and GitHub. In contrast, its performance drops in more intricate applications like Microsoft ToDo and XR software. As for VideoLLMs, their significantly poorer performance is attributed to two main factors: their inability to accurately interpret GUI content from user inputs and a lack of sufficient GUI-oriented pretraining, which is evident from their inadequate performance in basic captioning and description tasks. See Appendix D for BLEU and BERTScore, as well as detailed performance for complex tasks.
| Models | Setting | Caption | Complex Tasks | Conversation | Average | |||||
| Concise | Detailed | Static | Dyn. | Pred. | Round 1 | Round 2 | ||||
| ImageLLMs | Gemini-Pro-1.5 | R. | 3.659 | 2.837 | 2.969 | 2.822 | 3.450 | 3.608 | 3.845 | 3.339 |
| E. | 3.350 | 2.468 | 2.741 | 2.431 | 3.292 | 3.458 | 3.837 | 3.152 | ||
| Qwen-VL-Max | R. | 2.381 | 1.758 | 2.277 | 2.144 | 2.724 | 3.125 | 3.317 | 2.676 | |
| E. | 2.459 | 1.693 | 2.143 | 1.954 | 2.742 | 3.174 | 3.298 | 2.624 | ||
| H. | 2.474 | 1.711 | 2.137 | 2.032 | 2.834 | 3.223 | 3.257 | 2.651 | ||
| GPT-4V | R. | 3.579 | 2.676 | 3.243 | 3.011 | 3.630 | 3.925 | 4.131 | 3.589 | |
| E. | 3.141 | 2.301 | 2.927 | 2.627 | 3.541 | 3.844 | 4.103 | 3.407 | ||
| H. | 3.352 | 2.509 | 3.053 | 2.849 | 3.609 | 3.928 | 4.163 | 3.520 | ||
| C.C. | 3.454 | 2.547 | 1.818 | 2.335 | 3.577 | 3.521 | 3.884 | 3.028 | ||
| D.C. | 3.412 | 2.627 | 2.603 | 2.591 | 3.723 | 3.759 | 4.072 | 3.350 | ||
| H.+D.C. | 3.436 | 2.677 | 2.927 | 2.750 | 3.791 | 3.857 | 4.148 | 3.494 | ||
| GPT-4o | H. | 4.048 | 3.028 | 3.125 | 3.117 | 3.562 | 4.129 | 4.318 | 3.644 | |
| VideoLLMs | ChatUnivi | - | 1.587 | 1.240 | 1.705 | 1.656 | 2.524 | 2.698 | 3.366 | 2.389 |
| Minigpt4Video | - | 1.246 | 1.073 | 1.249 | 1.235 | 1.675 | 1.494 | 1.719 | 1.475 | |
| VideoChat2 | - | 1.992 | 1.312 | 1.812 | 1.682 | 2.158 | 2.342 | 2.720 | 2.144 | |
| GUI-Vid | - | 3.562 | 2.058 | 2.376 | 2.090 | 3.435 | 3.080 | 3.260 | 2.847 | |
Performance Variate in Different GUI Scenarios and Applications.
GPT-4V (Figure 7) and Gemini (Figure 16) excel in common scenarios such as mobile and website interfaces but show marked deficiencies in more complex GUI environments like XR and multi-window interactions, across both captioning and intricate tasks. This performance gap highlights a significant shortfall in understanding environments where GUI elements are scattered and demand sophisticated interpretation. It emphasizes the critical need for specialized benchmarks and datasets tailored to these complex GUI scenarios, which is essential for enhancing the GUI-oriented capabilities of MLLMs, paving the way for them to become truly reliable and high-performing general control agents.
Keyframe Selection is Important for GUI-oriented Tasks.
Across both basic tasks such as captioning and more complex tasks like prediction and reasoning, significant variations are evident among keyframe selection methods. As shown in Table 14 and Table 16, GPT-4V and Gemini significantly benefit from using random-selected and human-selected keyframes, scoring approximately 0.2-0.3 points higher in both captioning and free-form tasks than those using programmatic extraction. This suggests that traditional keyframe technologies, designed for natural videos, are less effective for detecting essential GUI operations, particularly when subtle movements like mouse clicks and dynamic changes are involved. Conversely, the difference in performance is relatively smaller in Qwen-VL-Max, indicating that while keyframe selection methods are crucial for models proficient in GUI content, they exert less influence on less capable models.
Dynamic GUI Tasks Continue to Challenge MLLMs.
In the fine-grained tasks depicted in Table 5, GPT-4V and GPT-4o excel with static GUI content and prediction tasks over image sequences but struggle with providing detailed descriptions for entire videos and dynamic GUI content. This discrepancy is attributed to minor variations in GUI that significantly impact descriptions. Enhancing the number of keyframes and the granularity of perception might mitigate these issues. Among VideoLLMs, ChatUnivi excels in conversational tasks by effectively leveraging contextual nuances, particularly in subsequent rounds, yet it underperforms in GUI-oriented captioning tasks. In contrast, GUI-Vid demonstrates proficiency in sequential tasks but falls short in both captioning and static content handling. This gap is linked to deficiencies in GUI-Vid’s pretraining, which lacked comprehensive GUI content crucial for effective vision-text alignment, as evidenced by its poor performance in Table 13 and an instruction tuning process also failed to fully address these shortcomings.
Vision Perception is Important for Sequential GUI Tasks.
As demonstrated in Table 5, integrating detailed textual information slightly outperforms purely vision-based inputs or detailed captions, akin to a Chain of Thought (CoT) [43] setting. Surprisingly, GPT-4V excels in caption and prediction tasks with just detailed captions, providing insights on enhancing specific GUI-oriented tasks through additional textual information. However, it still falls short in more challenging tasks, such as retrieving static or dynamic content. This underscores the critical role of visual perception in GUI environments, where even minor changes can significantly impact outcomes.
Supreme Enhancement of GUI-Vid on Graphic-based Interface After Finetuned on GUI-World.
As a pioneering study in training VideoLLMs as screen agents, GUI-Vid significantly outperforms the baseline model, showing an average improvement of 30% across various tasks and GUI scenarios, even surpassing the commercial ImageLLM, Qwen-VL-Max. This enhancement is particularly notable in captioning and prediction over image sequences, where GUI-Vid matches the performance of GPT-4V and Gemini-Pro. As shown in Figure 8, our two-stage progressive fintuning significantly enhances the performance in all GUI scenarios. Remarkably, GUI-Vid scored 3.747 in caption tasks within the XR scenario, highlighting its potential in XR applications and the high-quality annotations provided by our dataset. However, in Multiple-Choice QA and Chatbot tasks, GUI-Vid still lags behind industry leaders like GPT-4V and Gemini-Pro, a discrepancy likely due to the baseline LLM’s weaker performance and the challenges of instruction-based fine-tuning.
| Setting | F.K. | E.K. | Data | Software | Website | XR | Multi | IOS | Android | Avg. | ||||||||
| I. | V. | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |||
| Baseline | - | 8 | - | - | 45.5% | 2.144 | 42.6% | 2.221 | 44.0% | 2.005 | 40.4% | 2.222 | 40.2% | 2.169 | 44.7% | 2.119 | 42.9% | 2.147 |
| - | 16 | - | - | 45.1% | 2.144 | 41.8% | 2.240 | 41.0% | 2.007 | 40.7% | 2.238 | 39.9% | 2.138 | 44.7% | 2.147 | 42.2% | 2.154 | |
| GUI-Vid | 8 | 8 | ✘ | ✔ | 58.3% | 2.709 | 53.6% | 2.817 | 62.2% | 2.626 | 54.2% | 2.627 | 53.1% | 2.708 | 54.9% | 2.501 | 56.0% | 2.665 |
| ✔ | ✔ | 59.9% | 2.856 | 54.1% | 2.925 | 59.0% | 2.751 | 52.1% | 2.837 | 50.0% | 2.756 | 54.0% | 2.571 | 54.8% | 2.782 | |||
| 16 | ✘ | ✔ | 59.0% | 2.709 | 55.1% | 2.821 | 62.8% | 2.645 | 53.3% | 2.624 | 55.5% | 2.727 | 55.7% | 2.501 | 56.9% | 2.671 | ||
| ✔ | ✔ | 59.9% | 2.847 | 54.1% | 2.957 | 55.6% | 2.764 | 52.9% | 2.861 | 51.8% | 2.772 | 53.4% | 2.572 | 54.6% | 2.796 | |||
| Res. | Desc. | Conv. | Dyn. | Static | Caption | Average |
| Low | 2.794 | 3.912 | 3.150 | 2.869 | 3.672 | 3.394 |
| High | 3.031 | 4.056 | 3.318 | 3.131 | 3.911 | 3.573 |
Upper Bound of GUI-oriented Capability with More Keyframes and High Resolution.
As depicted in Table 6, our two ablation studies during the fine-tuning phase demonstrate that utilizing GUI image-text captioning data significantly enhances the model’s preliminary understanding of GUI elements, outperforming training that relies solely on videos. Additionally, an increased number of keyframes correlates with improved performance across various scenarios, notably in environments featuring multiple windows and software applications. Further evidence from Table 7 reveals that higher image resolutions substantially boost task performance, both basic and complex, for GPT-4o. These findings underscore the potential for further developing a more robust GUI Agent.
5 Related Work
MLLM-based Agents for GUI.
Building upon the significant advancements in LLMs [52, 53, 54, 55] and advanced modality-mixing technologies [56, 57], groundbreaking MLLMs such as GPT-4V [1] and Gemini-Pro [42], along with open-source MLLMs like the LLaVA-1.6 series [2, 58], CogVLM [59], and Qwen-VL series [41], have shown outstanding performance across various tasks [60, 61, 62, 63, 64, 65, 66, 67, 68]. Venturing beyond text and single image, several studies are now exploring the integration of video modalities for tasks requiring dynamic or sequential visual content [44, 35, 69, 70]. In the GUI domain, leveraging the robust vision perception capabilities of MLLMs, applications such as WebAgents [8, 71, 23] and Mobile Agents [17, 12, 72] have gained popularity for handling everyday tasks like navigation and VQA. Frontier research is also investigating the use of MLLMs as general control agents, such as in playing computer games [73, 74] and serving as OS co-pilots [75, 24], paving the way for more complex GUI operations.
GUI Benchmark & Dataset.
Building upon the foundational work of Rico [13], the first mobile GUI video dataset, and AitW [16], which features 715k episodes of sequential images, research has extensively covered mobile [14, 76, 77] and web GUI environments [78, 19, 79, 80, 81]. Mind2Web [20] stands out in web-based datasets with over 2,000 tasks from 137 websites across 31 domains. Advances continue into desktop GUIs with new toolkits [23], benchmarks [21, 82], and frameworks [83, 84, 11]. Research on GUI also transfers from comprehending single images in a static workspace [8] to sequential operations or multi-hop scenarios [24, 22], challenging the understanding and operation capability of these powerful models.
6 Conclusion
In this paper, we have introduced GUI-World, a comprehensive GUI-oriented dataset designed to benchmark and enhance understanding of virtual interfaces, especially sequential and dynamic tasks. This dataset extensively covers six scenarios and various tasks, addressing the previous research gap in comprehensively evaluating models’ capabilities in graphic-based understanding. We conduct extensive benchmarks on leading MLLMs and the first Video Agent ‘GUI-Vid’ finetuned on our GUI-World specifically for tasks requiring temporal information, achieving results comparable to top-performing models, providing detailed insights into enhancing GUI-related capabilities.
7 Limitations
While our work presents significant advancements in the field of GUI agents, there are several limitations that need to be addressed. Firstly, despite expanding the dataset to include various GUI scenarios, our models still show limited generalization capabilities when applied to environments not represented in the training data. This highlights the need for further research to improve the adaptability and robustness of GUI agents in diverse and unseen environments. Additionally, the accuracy of our models heavily relies on the selection of keyframes. Automatically extracted keyframes often fail to capture the essential elements needed for accurate GUI understanding, indicating the need for more sophisticated keyframe extraction techniques. Furthermore, although VideoLLMs have shown improvements in handling dynamic content, their ability to understand and predict sequential information in GUI tasks remains suboptimal. This suggests a necessity for future work to focus on enhancing the temporal understanding capabilities of these models. Finally, the training and fine-tuning processes for VideoLLMs require significant computational resources, which may not be accessible to all researchers.
8 Potential Negative Societal Impacts
While our work aims to advance the capabilities of GUI agents for beneficial applications, it is important to consider potential negative societal impacts. The use of GUI agents, especially those capable of operating across multiple environments and platforms, raises significant privacy concerns. Ensuring that these agents operate within strict ethical guidelines and that user data is handled securely and responsibly is paramount. There is also the risk of misuse of advanced GUI agents for malicious purposes, such as unauthorized access to sensitive information or automated exploitation of software vulnerabilities. Establishing robust security measures and ethical usage policies is essential to mitigate these risks.
Appendix A Details of Dataset Construction
A.1 Six Main GUI Categories
In earlier endeavors pertaining to GUI, such as those involving GUI testing [85, 86, 87], the focus was segmented into GUIs for Website, Software, IOS and Android platforms. However, as a comprehensive GUI dataset, we included all potential GUI scenarios in our dataset to ensure that our data is the most comprehensive knowledge that the GUI Agent needs to learn; we divided these scenarios into six categories:
Android. This category focuses on the GUI scenarios that occur within the Android operating system, which is predominantly used on smartphones. Android’s ubiquity in the mobile market has led to a wide variety of GUI designs and interaction patterns, making it a rich field for study. This category has been the subject of extensive scrutiny in scholarly works such as [13, 76, 16, 88].
Software. This category encapsulates the GUI scenarios arising within software applications, whether they are standalone programs or components of a larger suite. The diversity of software applications, from productivity tools to creative suites, offers a wide range of GUI scenarios for exploration. The literature is rich with research in this area, such as [89].
Website. This category is concerned with the GUI scenarios that manifest within a web browser. Given the ubiquity of web browsing in modern digital life, this category holds significant relevance. It holds a substantial representation in academic literature, with pioneering papers such as [20, 21] proposing excellent GUI datasets for websites.
IOS. This category zeroes in on the GUI scenarios that transpire within the iOS operating system, the proprietary system for Apple devices like the iPhone and iPad. The iOS platform is known for its distinct design aesthetics and interaction patterns, providing a unique context for GUI research. A number of studies, such as [90, 91] make use of GUI information in IOS.
Multi Windows. This category is dedicated to GUI scenarios that necessitate simultaneous interaction with multiple windows, a common occurrence in desktop environments where users often juggle between several applications or documents. Despite the common use of multi-window interaction in everyday GUI usage, there has been relatively little research into this area [92]. The need for efficient multitasking in such scenarios presents unique challenges and opportunities for GUI design and interaction research. As of our knowledge, there are no specific datasets catering to these multi-window GUI scenarios.
XR. XR encompasses Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR) [93]. Given the advancements in XR technology and the growing accessibility of commercial-grade head-mounted displays [25, 94], XR has emerged as a novel medium for human-computer interaction. This necessitates the exploration of GUI within XR environments. In these scenarios, the GUI takes on a 3D, immersive form [95], demanding the agent to comprehend and navigate a 3D space. The emerging field of XR presents a new frontier for GUI research, with unique challenges and opportunities due to its immersive and interactive nature. To date, as far as we are aware, there are no datasets that specifically address GUI in the realm of XR.
A.2 Selected Website/Software
In our study, we selected a diverse range of websites and software to comprehensively evaluate GUI understanding capabilities across various user scenarios. These selections cover essential categories such as social media, productivity tools, online shopping, and educational platforms, providing a broad spectrum of GUI environments.
The chosen websites, as shown in Figure 9, include popular social media platforms like Instagram, Twitter, and LinkedIn, which are integral to understanding dynamic and interactive GUI elements. We also included widely-used productivity tools such as Microsoft Teams, Notion, and Slack to evaluate GUI tasks in professional and collaborative settings.
For software shown in Figure 10, we incorporated key applications like Adobe Photoshop and MATLAB to assess GUI operations in specialized and technical environments. Additionally, video conferencing tools like Zoom and cloud storage services like Google Drive were included to represent common remote work and file management scenarios.
These selections ensure that our study encompasses a wide array of user interactions and GUI complexities, thereby providing a robust evaluation of the current state-of-the-art methods in GUI understanding by MLLMs and comprehensively constructing a high-quality dataset.
A.3 Human Keyframes Annotation Process
Annotator’s Information
The annotation is conducted by 16 authors of this paper and 8 volunteers independently. As acknowledged, the diversity of annotators plays a crucial role in reducing bias and enhancing the reliability of the benchmark. These annotators have knowledge in the GUI domain, with different genders, ages, and educational backgrounds. The education backgrounds of annotators are above undergraduate. To ensure the annotators can proficiently mark the data, we provide them with detailed tutorials, teaching them how to use software to record videos or edit video clips. We also provide them with detailed criteria and task requirements in each annotation process.
Recording Video.
For self-recording videos, we employ OBS333https://obsproject.com/ on the Windows system for screen capturing and the official screen recording toolkit on the Mac/IOS system. This process necessitates human labelers to execute a series of targeted actions within specific websites or applications, which are subsequently captured as raw video footage. These actions, commonplace in everyday usage, enhance the reliability of our dataset. Subsequently, the raw videos are segmented into sub-videos, each encapsulating multiple actions (e.g., clicking a button) to achieve a specific objective (e.g., image search). The videos are then processed to extract keyframes annotated with detailed descriptions.
Edition Based on YouTube Videos.
For sourcing videos from YouTube, we utilize a search protocol formatted as "[website name/application name] + tutorial" to compile relevant video lists. Human labelers first review these videos to understand the primary operations they depict. These videos are then divided into sub-videos, each containing several actions directed towards a single goal (e.g., image search). Like the self-recorded footage, these segments are processed to isolate keyframes and furnish them with descriptive annotations.
Keyframes Annotation.
After obtaining the GUI video clips, human annotators will filter out the keyframes of the operations based on the video content and the mouse and keyboard actions at that time. They will also label the sub-operations or targets between the two keyframes. Once the annotation is complete, the annotators will provide an overall description of the entire video, summarizing the main goal of the human operations in the video. After all the information is annotated, we will use a Large Language Model (LLM) to refine the text content, reducing any errors made by human annotators and adjusting the sentence structure. The prompt we use for the LLM to polish the human annotations is shown in Figure 11 and Figure 12.
Human-LLM Cooperated Instruction Generation.
To curate and refine the golden answer of each video-instruction pair generated by GPT-4V, given that the raw response from GPT-4V may contain harmful content or hallucinations. The role of humans in the golden answer generation process is to enhance the difficulty of the questions and remove harmful and incorrect content, as shown in Table 8.
| Raw response from GPT-4V | Golden answer from humans |
| "Q": "What subcategories are visible under the ’Lifestyle’ section in the static GUI content of the video?" | "Q": "What subcategories are visible under the ’Lifestyle’ section in the static GUI content of the video?" |
| "A": "In the ’Lifestyle’ section of the telegram.com website, the visible subcategories are ’Worcester Living’, ’Faith’, and ’Food’. These are clearly displayed in a horizontal list just below the main ’Lifestyle’ category header, indicating different areas of content that users can explore within the lifestyle section." | "A": "The ’Lifestyle’ section of the telegram.com website features visible subcategories such as ’Worcester Living’, ’Faith’, and ’Food’. These are prominently shown in a horizontal list beneath the main ’Lifestyle’ header, offering users various content areas to explore within the section." |
Human verifying GPT-4V annotated captions.
We evaluated the quality of annotations from GPT-4V by selecting 1,000 detailed descriptions and captions generated by GPT-4V, which were then assessed by human annotators. The high satisfaction rate of 98% underscores the quality and relevance of the GPT-4V annotations.
Appendix B Dataset Analysis
In this section, we provide an analysis of the length distribution of QA in each GUI scenario, as illustrated in Figure 13 and Figure 14. Questions focused on sequential and predictional tasks are slightly longer than other types, while the golden answer of static tasks tends to be longer. Length of Question-answer pair in various GUI scenarios is similarly distributed, with questions in Android environment being slightly shorter, and answers in XR environment being longer.
Appendix C Details of Experiments Setups
C.1 Finetune dataset construction
We use two settings to finetune GUI-Vid, one with video-text pairs only, and the other with video-text and image-text pairs, which are all GUI content:
Video Only. In this setting, we only trained GUI-Vid with video-text pairs in GUI-World, as shown in Table 9.
Video-Image. Inspired by the pre-trained process of Videochat2, we include image-text pairs to help the visual encoder align GUI knowledge. These images are selected from our GUI-World, MetaGUI [14], and OmniAct [21] for high-quality GUI content. Subsequently, we use GPT-4V to generate a detailed description and a concise caption for each image. Finally, we construct a dataset consisting of video-text and image-text pairs for gaining comprehensive GUI-oriented capabilities.
| Stage | Data types | Amount |
| 1 | Detailed Description | 14,276 |
| Concise Caption | 7,138 | |
| 2 | GUI VQA | 21,414 |
| Multiple-Choice QA | 14,276 | |
| Conversation | 7,138 |
| Stage | Data types | Source | Type | Amount |
| 1 | GUI-World | Video | Detailed Description | 14,276 |
| Concise Caption | 7,138 | |||
| Image | Detailed Description | 5,555 | ||
| Concise Caption | 5,555 | |||
| MetaGUI | Image | Detailed Description | 19,626 | |
| Concise Caption | 19,626 | |||
| OmniAct | Detailed Description | 260 | ||
| Concise Caption | 260 | |||
| 2 | GUI-World | Video | GUI VQA | 21,414 |
| Multiple-Choice QA | 14,276 | |||
| Conversation | 7,138 |
C.2 Hyperparameter Settings
In this section, we will introduce the hyperparameters of MLLMs to facilitate experiment reproducibility and transparency. We divide them into three parts: the inference phase during benchmark and dataset construction, the LLM-as-a-Judge phase, and the fine-tuning phase. All our experiments were conducted on a server equipped with dual A800 and dual 4090 GPUs.
Inference.
We empirically study 7 MLLMs, involving 4 Image-LLMs and 3 Video-LLMs, with their hyperparameters detailed as follows:
Gemini-Pro-1.5 [42]: We use the default settings, which set temperature as 0.4, top-p as 1, and max-token as 2048. It should be noted that during our project, Gemini-Pro-1.5 is still under the user request limit, which only provides 100 requests per day, making our benchmark difficult. Given that Gemini hasn’t launched Pay-as-you-go444https://ai.google.dev/pricing, we will include benchmark results on ‘Human’ setting as soon as possible.
Qwen-VL-Max [41]: We use the default settings for Qwen-VL-Max, with top-p as 0.8 and max-token as 2048. Given that the input context window is merely 6,000 for Qwen, we scale the resolution for all images to 0.3.
ChatUnivi [44]: We use ChatUnivi-7B built upon Vicuna-v0-7B and set the max frame as 100, temperature as 0.2, and max-token as 1024.
Minigpt4video [45]: We use the suggested settings555https://github.com/Vision-CAIR/MiniGPT4-video for this model and the max-frame are set as 45, with only the max-token being modified to 1024.
VideoChat2 & GUI-Vid [46]: For a fair comparison, we set the same hyperparameters for VideoChat2 & GUI-Vid. We set the max-token as 1024, top-p as 0.9, temperature as 1.0, max-frame as 8/16, repetition penalty as 1.2, and length penalty as 1.2.
LLM-as-a-Judge.
We studied four LLM-as-a-Judge in giving a similarity score for the MLLM’s response and ground truth, namely GPT-4 [52], ChatGPT [96], LLaMA-3-70b-instruct [54], and Mixtral-8x22b-instruct-v0.1 [55]. Hyperparameter settings are detailed as follows:
GPT-4 & ChatGPT. We set the temperature as 0.6 and others as default.
LLaMA-3-70b-instruct. We set the temperature as 0.6, top-p as 0.9, top-k as 50.
Mixtral-8x22b-instruct-v0.1. We set top-p as 0.7, top-k as 50, and temperature as 0.7.
Finetune.
We include several hyperparameter settings in experiment settings and ablation studies, as shown in Table 11.
| Config | Setting |
| input frame | 8 |
| input resolution | 224 |
| max text length | 512 |
| input modal | I. + V. |
| optimizer | AdamW |
| optimizer momentum | |
| weight decay | 0.02 |
| learning rate schedule | cosine decay |
| learning rate | 2e-5 |
| batch size | 4 |
| warmup epochs | 0.6 |
| total epochs | 3 |
| backbone drop path | 0 |
| QFormer drop path | 0.1 |
| QFormer dropout | 0.1 |
| QFormer token | 96 |
| flip augmentation | yes |
| augmentation | MultiScaleCrop [0.5, 1] |
| Models | Pearson() | Spearman() | Kendall() | $ per Benchmark() |
| GPT-4 | 0.856 | 0.853 | 0.793 | $ |
| ChatGPT | 0.706 | 0.714 | 0.627 | 12$ |
| Llama-3-70b-instruct | 0.774 | 0.772 | 0.684 | 12$ |
| Mixtral-8x22b-instruct-v0.1 | 0.759 | 0.760 | 0.670 | $ |
C.3 Evaluation.
Given the complexity of free-form answers in GUI scenarios, the evaluation includes specific positions of GUI elements, textual content, and comparing the response to the golden answer. LLM-as-a-judge has been widely used in previous studies for complex evaluation tasks [47, 48]. Therefore, we leverage LLM-as-a-Judge [47] in a similar setting to MM-vet [60], which compares the MLLM’s response to the golden answer. We carefully evaluate the accessibility of leveraging LLM-as-a-Judge, selecting 1,000 samples covering 6 free-form questions mentioned in our dataset. As shown in Table 12, GPT-4 outperforms other LLMs, exhibiting a better human alignment on providing a similarity score for the response compared to the golden answer, although it is approximately 10 times more expensive than other models.
Appendix D Additional Experiments Results
In this section, we provide detailed results on each task in each GUI scenario. For captioning tasks, Table 13 shows comprehensive experimental results among six scenarios. For scores of LLM-as-a-Judge in a specific task, see Table 14, Table 15, Table 16, Table 17, and Table 18. For BLEU [50] and BERTScore [51] in validating free-form and conversational questions, see Table 19, Table 20, Table 21, Table 24, Table 22, and Table 23. For performance in fine-grain (application level), see Figure 16 for Gemini-Pro and Figure 17 for Qwen-VL-Max.
| Models | Setting | Software | Website | XR | Multi | IOS | Android | Avg. | |||||||
| Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | ||
| Gemini-Pro-1.5 | R. | 3.659 | 2.837 | 3.613 | 2.860 | 2.995 | 2.590 | 3.276 | 2.470 | 3.678 | 2.936 | - | - | 3.444 | 2.739 |
| E. | 3.350 | 2.468 | 3.159 | 2.422 | 2.837 | 2.279 | 2.824 | 2.109 | 3.394 | 2.519 | 3.185 | 2.312 | 3.125 | 2.351 | |
| Qwen-VL-Max | R. | 2.381 | 1.758 | 2.326 | 1.681 | 2.172 | 1.772 | 2.035 | 1.463 | 2.513 | 1.662 | 2.141 | 1.565 | 2.261 | 1.650 |
| E. | 2.459 | 1.693 | 2.317 | 1.599 | 2.167 | 1.638 | 2.190 | 1.438 | 2.189 | 1.615 | 2.002 | 1.429 | 2.221 | 1.569 | |
| H. | 2.474 | 1.711 | 2.457 | 1.698 | 2.383 | 1.777 | 1.910 | 1.346 | 2.577 | 1.795 | 2.474 | 1.711 | 2.360 | 1.665 | |
| GPT-4V | R. | 3.579 | 2.676 | 3.612 | 2.699 | 2.975 | 2.525 | 3.281 | 2.661 | 3.757 | 2.775 | 3.655 | 2.755 | 3.479 | 2.682 |
| E. | 3.141 | 2.301 | 3.293 | 2.380 | 2.471 | 2.085 | 3.063 | 2.324 | 3.624 | 2.611 | 3.201 | 2.312 | 3.132 | 2.335 | |
| H. | 3.352 | 2.509 | 3.702 | 2.750 | 3.050 | 3.556 | 3.524 | 2.673 | 3.670 | 2.588 | - | - | 3.460 | 2.614 | |
| GPT-4o | H. | 4.048 | 3.028 | 4.067 | 3.233 | 3.398 | 2.729 | 3.869 | 3.111 | 4.014 | 2.993 | 4.071 | 3.095 | 3.911 | 3.869 |
| ChatUnivi | - | 1.587 | 1.240 | 1.569 | 1.254 | 1.417 | 1.148 | 1.575 | 1.267 | 1.480 | 1.146 | 1.778 | 1.249 | 1.568 | 1.217 |
| Minigpt4Video | - | 1.246 | 1.073 | 1.200 | 1.057 | 1.320 | 1.106 | 1.130 | 1.034 | 1.190 | 1.076 | 1.184 | 1.061 | 1.212 | 1.068 |
| VideoChat2 | - | 1.992 | 1.312 | 1.817 | 1.307 | 1.838 | 1.426 | 2.222 | 1.433 | 2.169 | 1.270 | 2.119 | 1.294 | 1.900 | 1.340 |
| GUI-Vid | - | 3.562 | 2.085 | 3.655 | 2.167 | 3.747 | 2.153 | 3.370 | 1.742 | 3.566 | 2.071 | 2.662 | 1.248 | 3.427 | 1.911 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 3.279 | 3.050 | 3.560 | 3.579 | 3.796 | 3.452 |
| E. | 2.983 | 2.491 | 3.432 | 3.405 | 3.760 | 3.215 | |
| Qwen-VL-Max | R. | 2.317 | 2.271 | 2.802 | 2.995 | 3.069 | 2.656 |
| E. | 2.256 | 2.198 | 2.821 | 2.861 | 3.144 | 2.627 | |
| H. | 2.308 | 2.078 | 2.832 | 3.061 | 3.358 | 2.698 | |
| GPT-4V | R. | 3.461 | 3.214 | 3.754 | 3.778 | 4.029 | 3.648 |
| E. | 3.197 | 2.808 | 3.487 | 3.717 | 3.954 | 3.433 | |
| H. | 3.498 | 3.255 | 3.727 | 3.731 | 4.061 | 3.655 | |
| C.C. | 1.746 | 2.738 | 3.645 | 3.363 | 3.632 | 3.025 | |
| D.C. | 2.704 | 2.917 | 3.686 | 3.680 | 3.901 | 3.380 | |
| H.+D.C. | 3.313 | 3.221 | 3.852 | 3.850 | 4.171 | 3.682 | |
| GPT-4o | H. | 3.443 | 3.373 | 3.672 | 4.086 | 4.122 | 3.740 |
| ChatUnivi | - | 1.701 | 1.668 | 2.524 | 2.514 | 3.338 | 2.349 |
| Minigpt4Video | - | 1.309 | 1.233 | 1.766 | 1.439 | 1.854 | 1.520 |
| VideoChat2 | - | 1.771 | 1.777 | 2.288 | 2.461 | 2.812 | 2.221 |
| GUI-Vid | - | 2.406 | 2.341 | 3.544 | 3.135 | 3.355 | 2.957 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 2.892 | 2.505 | 3.543 | 3.222 | 3.611 | 3.154 |
| E. | 2.814 | 2.163 | 3.510 | 3.108 | 3.455 | 3.006 | |
| Qwen-VL-Max | R. | 2.047 | 1.968 | 2.712 | 2.879 | 3.132 | 2.469 |
| E. | 2.125 | 1.973 | 2.658 | 2.760 | 3.029 | 2.499 | |
| H. | 1.886 | 1.920 | 2.656 | 2.727 | 3.012 | 2.373 | |
| GPT-4V | R. | 2.934 | 2.668 | 3.392 | 3.291 | 3.714 | 3.200 |
| E. | 2.222 | 2.153 | 3.310 | 3.151 | 3.618 | 2.892 | |
| H. | 2.893 | 2.778 | 3.538 | 3.364 | 3.747 | 3.265 | |
| C.C. | 1.744 | 2.412 | 3.327 | 3.080 | 3.485 | 2.809 | |
| D.C. | 2.427 | 2.409 | 3.518 | 3.176 | 3.749 | 3.056 | |
| H.+D.C. | 2.775 | 2.635 | 3.580 | 3.235 | 3.734 | 3.191 | |
| GPT-4o | H. | 2.871 | 2.745 | 3.370 | 3.596 | 3.836 | 3.285 |
| ChatUnivi | - | 1.660 | 1.420 | 2.205 | 2.250 | 3.270 | 2.161 |
| Minigpt4Video | - | 1.225 | 1.161 | 1.610 | 1.347 | 1.465 | 1.362 |
| VideoChat2 | - | 1.654 | 1.547 | 2.192 | 2.099 | 2.529 | 2.005 |
| GUI-Vid | - | 2.444 | 2.147 | 3.347 | 2.836 | 3.036 | 2.764 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 2.538 | 2.410 | 3.296 | 3.152 | 3.402 | 2.959 |
| E. | 2.545 | 2.049 | 2.972 | 2.930 | 3.389 | 2.777 | |
| Qwen-VL-Max | R. | 1.793 | 1.872 | 2.770 | 2.897 | 3.122 | 2.432 |
| E. | 1.866 | 1.780 | 2.730 | 2.627 | 3.105 | 2.362 | |
| H. | 1.884 | 1.969 | 2.913 | 2.689 | 3.104 | 2.490 | |
| GPT-4V | R. | 3.185 | 2.655 | 3.745 | 3.699 | 3.973 | 3.452 |
| E. | 2.902 | 2.406 | 3.636 | 3.420 | 3.729 | 3.219 | |
| H. | 3.000 | 2.952 | 3.801 | 3.597 | 3.889 | 3.449 | |
| C.C. | 2.097 | 2.973 | 3.774 | 3.331 | 3.621 | 3.160 | |
| D.C. | 2.671 | 2.979 | 3.849 | 3.466 | 3.822 | 3.358 | |
| H.+D.C. | 3.037 | 3.162 | 4.079 | 3.748 | 4.036 | 3.617 | |
| GPT-4o | H. | 3.108 | 3.106 | 3.829 | 4.043 | 4.188 | 3.654 |
| ChatUnivi | - | 1.658 | 1.623 | 2.514 | 2.384 | 3.199 | 2.275 |
| Minigpt4Video | - | 1.205 | 1.186 | 1.690 | 1.400 | 1.801 | 1.457 |
| VideoChat2 | - | 1.754 | 1.774 | 2.479 | 2.420 | 2.699 | 2.222 |
| GUI-Vid | - | 2.485 | 2.067 | 3.537 | 2.954 | 3.247 | 2.861 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 3.076 | 2.637 | 3.370 | 3.366 | 3.615 | 3.213 |
| E. | 2.852 | 2.356 | 3.137 | 3.126 | 3.566 | 3.007 | |
| Qwen-VL-Max | R. | 2.438 | 2.244 | 2.923 | 3.102 | 3.273 | 2.779 |
| E. | 2.303 | 2.150 | 2.614 | 3.145 | 3.264 | 2.659 | |
| H. | 1.884 | 1.969 | 2.913 | 2.689 | 3.104 | 2.490 | |
| GPT-4V | R. | 3.364 | 3.080 | 3.684 | 3.766 | 4.184 | 3.614 |
| E. | 3.209 | 2.774 | 3.545 | 3.611 | 4.006 | 3.427 | |
| H. | 3.107 | 2.830 | 3.631 | 3.680 | 4.011 | 3.453 | |
| C.C. | 1.788 | 2.291 | 3.511 | 3.212 | 3.542 | 2.868 | |
| D.C. | 2.751 | 2.732 | 3.654 | 3.642 | 3.842 | 3.324 | |
| H.+D.C. | 3.090 | 2.965 | 3.740 | 3.786 | 3.994 | 3.516 | |
| GPT-4o | H. | 3.183 | 2.993 | 3.460 | 4.050 | 4.141 | 3.558 |
| ChatUnivi | - | 1.771 | 1.642 | 2.408 | 2.559 | 3.307 | 2.337 |
| Minigpt4Video | - | 1.291 | 1.219 | 1.698 | 1.556 | 1.737 | 1.501 |
| VideoChat2 | - | 1.955 | 1.803 | 2.145 | 2.315 | 2.626 | 2.169 |
| GUI-Vid | - | 2.262 | 2.133 | 3.401 | 2.843 | 3.224 | 2.773 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | E. | 2.703 | 2.460 | 3.157 | 3.642 | 3.881 | 3.168 |
| Qwen-VL-Max | R. | 1.887 | 1.804 | 2.398 | 2.823 | 3.056 | 2.309 |
| E. | 1.785 | 1.630 | 2.311 | 2.605 | 3.233 | 2.277 | |
| GPT-4V | R. | 3.116 | 3.047 | 3.477 | 3.924 | 4.008 | 3.515 |
| E. | 2.705 | 2.470 | 3.175 | 3.647 | 3.885 | 3.176 | |
| C.C. | 2.092 | 2.243 | 3.139 | 3.443 | 3.782 | 2.939 | |
| D.C. | 3.015 | 2.890 | 3.357 | 3.883 | 3.990 | 3.427 | |
| GPT-4o | H. | 3.057 | 3.220 | 3.373 | 3.981 | 4.186 | 3.561 |
| ChatUnivi | - | 1.835 | 1.654 | 2.317 | 2.712 | 3.433 | 2.390 |
| Minigpt4Video | - | 1.183 | 1.159 | 1.507 | 1.342 | 1.521 | 1.342 |
| VideoChat2 | - | 1.732 | 1.754 | 2.125 | 2.340 | 2.645 | 2.119 |
| GUI-Vid | - | 2.010 | 1.928 | 3.053 | 2.755 | 3.105 | 2.572 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | R. | 0.109 | 0.789 | 0.150 | 0.720 | 0.078 | 0.680 | 0.056 | 0.716 | 0.016 | 0.605 | 0.122 | 0.761 | 0.089 | 0.712 |
| E. | 0.093 | 0.758 | 0.134 | 0.699 | 0.072 | 0.659 | 0.046 | 0.682 | 0.011 | 0.558 | 0.106 | 0.747 | 0.077 | 0.684 | |
| Qwen-VL-Max | R. | 0.085 | 0.698 | 0.101 | 0.649 | 0.064 | 0.576 | 0.010 | 0.521 | 0.008 | 0.443 | 0.121 | 0.749 | 0.065 | 0.606 |
| E. | 0.094 | 0.704 | 0.103 | 0.633 | 0.062 | 0.595 | 0.009 | 0.524 | 0.006 | 0.437 | 0.113 | 0.739 | 0.065 | 0.605 | |
| H. | 0.081 | 0.676 | 0.098 | 0.620 | 0.067 | 0.596 | 0.009 | 0.504 | 0.004 | 0.429 | 0.117 | 0.743 | 0.063 | 0.595 | |
| GPT-4V | R. | 0.162 | 0.814 | 0.206 | 0.753 | 0.190 | 0.739 | 0.041 | 0.676 | 0.033 | 0.581 | 0.181 | 0.793 | 0.136 | 0.726 |
| E. | 0.161 | 0.792 | 0.191 | 0.726 | 0.175 | 0.724 | 0.030 | 0.609 | 0.017 | 0.486 | 0.165 | 0.786 | 0.123 | 0.687 | |
| H. | 0.153 | 0.805 | 0.194 | 0.737 | 0.183 | 0.731 | 0.037 | 0.639 | 0.025 | 0.537 | 0.179 | 0.791 | 0.129 | 0.707 | |
| GPT-4o | H. | 0.131 | 0.806 | 0.212 | 0.776 | 0.147 | 0.728 | 0.041 | 0.711 | 0.018 | 0.575 | 0.159 | 0.803 | 0.118 | 0.733 |
| ChatUnivi | - | 0.097 | 0.697 | 0.074 | 0.581 | 0.101 | 0.619 | 0.005 | 0.409 | 0.000 | 0.195 | 0.084 | 0.723 | 0.060 | 0.537 |
| Minigpt4Video | - | 0.019 | 0.516 | 0.022 | 0.470 | 0.029 | 0.516 | 0.000 | 0.399 | 0.000 | 0.249 | 0.013 | 0.510 | 0.014 | 0.443 |
| VideoChat2 | - | 0.095 | 0.698 | 0.080 | 0.595 | 0.076 | 0.574 | 0.004 | 0.341 | 0.000 | 0.193 | 0.100 | 0.733 | 0.059 | 0.523 |
| GUI-Vid | - | 0.142 | 0.758 | 0.145 | 0.681 | 0.114 | 0.698 | 0.049 | 0.658 | 0.004 | 0.519 | 0.093 | 0.717 | 0.091 | 0.672 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | R. | 0.113 | 0.793 | 0.145 | 0.727 | 0.083 | 0.676 | 0.054 | 0.720 | 0.016 | 0.664 | 0.098 | 0.736 | 0.085 | 0.719 |
| E. | 0.095 | 0.754 | 0.121 | 0.681 | 0.079 | 0.661 | 0.041 | 0.676 | 0.011 | 0.602 | 0.092 | 0.725 | 0.073 | 0.683 | |
| Qwen-VL-Max | R. | 0.099 | 0.728 | 0.099 | 0.634 | 0.080 | 0.610 | 0.008 | 0.519 | 0.005 | 0.471 | 0.085 | 0.694 | 0.063 | 0.609 |
| E. | 0.083 | 0.710 | 0.101 | 0.631 | 0.093 | 0.611 | 0.011 | 0.503 | 0.004 | 0.469 | 0.099 | 0.709 | 0.065 | 0.605 | |
| H. | 0.079 | 0.693 | 0.089 | 0.597 | 0.093 | 0.606 | 0.009 | 0.488 | 0.007 | 0.449 | 0.103 | 0.705 | 0.063 | 0.590 | |
| GPT-4V | R. | 0.173 | 0.830 | 0.241 | 0.765 | 0.205 | 0.751 | 0.040 | 0.694 | 0.032 | 0.645 | 0.164 | 0.763 | 0.142 | 0.741 |
| E. | 0.159 | 0.802 | 0.204 | 0.727 | 0.202 | 0.727 | 0.033 | 0.648 | 0.031 | 0.590 | 0.149 | 0.757 | 0.130 | 0.708 | |
| H. | 0.182 | 0.823 | 0.234 | 0.771 | 0.213 | 0.758 | 0.043 | 0.696 | 0.041 | 0.660 | 0.165 | 0.768 | 0.147 | 0.746 | |
| GPT-4o | H. | 0.141 | 0.813 | 0.219 | 0.768 | 0.199 | 0.731 | 0.054 | 0.700 | 0.026 | 0.602 | 0.146 | 0.755 | 0.131 | 0.728 |
| ChatUnivi | - | 0.078 | 0.645 | 0.068 | 0.581 | 0.102 | 0.607 | 0.008 | 0.399 | 0.000 | 0.192 | 0.061 | 0.661 | 0.053 | 0.514 |
| Minigpt4Video | - | 0.022 | 0.527 | 0.016 | 0.448 | 0.027 | 0.501 | 0.000 | 0.344 | 0.000 | 0.186 | 0.011 | 0.522 | 0.013 | 0.421 |
| VideoChat2 | - | 0.073 | 0.619 | 0.075 | 0.579 | 0.049 | 0.511 | 0.004 | 0.328 | 0.000 | 0.167 | 0.067 | 0.678 | 0.045 | 0.480 |
| GUI-Vid | - | 0.114 | 0.731 | 0.158 | 0.674 | 0.129 | 0.694 | 0.049 | 0.667 | 0.002 | 0.553 | 0.075 | 0.681 | 0.088 | 0.667 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | R. | 0.088 | 0.772 | 0.101 | 0.678 | 0.070 | 0.678 | 0.026 | 0.650 | 0.002 | 0.463 | 0.082 | 0.733 | 0.062 | 0.662 |
| E. | 0.073 | 0.760 | 0.090 | 0.651 | 0.062 | 0.666 | 0.015 | 0.618 | 0.002 | 0.449 | 0.084 | 0.720 | 0.054 | 0.644 | |
| Qwen-VL-Max | R. | 0.069 | 0.703 | 0.075 | 0.602 | 0.049 | 0.601 | 0.006 | 0.486 | 0.000 | 0.338 | 0.117 | 0.738 | 0.053 | 0.578 |
| E. | 0.048 | 0.689 | 0.079 | 0.657 | 0.058 | 0.605 | 0.005 | 0.498 | 0.000 | 0.359 | 0.112 | 0.739 | 0.050 | 0.591 | |
| H. | 0.051 | 0.651 | 0.073 | 0.593 | 0.044 | 0.591 | 0.004 | 0.493 | 0.001 | 0.357 | 0.101 | 0.726 | 0.046 | 0.569 | |
| GPT-4V | R. | 0.093 | 0.794 | 0.169 | 0.715 | 0.165 | 0.736 | 0.028 | 0.625 | 0.006 | 0.457 | 0.147 | 0.768 | 0.101 | 0.683 |
| E. | 0.085 | 0.726 | 0.131 | 0.665 | 0.162 | 0.724 | 0.020 | 0.541 | 0.003 | 0.382 | 0.141 | 0.760 | 0.090 | 0.633 | |
| H. | 0.091 | 0.797 | 0.181 | 0.732 | 0.180 | 0.744 | 0.027 | 0.630 | 0.006 | 0.471 | 0.154 | 0.773 | 0.106 | 0.691 | |
| GPT-4o | H. | 0.077 | 0.800 | 0.154 | 0.717 | 0.153 | 0.718 | 0.020 | 0.615 | 0.006 | 0.468 | 0.138 | 0.759 | 0.091 | 0.680 |
| ChatUnivi | - | 0.083 | 0.686 | 0.061 | 0.538 | 0.091 | 0.575 | 0.006 | 0.475 | 0.000 | 0.282 | 0.086 | 0.693 | 0.054 | 0.541 |
| Minigpt4Video | - | 0.014 | 0.545 | 0.016 | 0.466 | 0.027 | 0.502 | 0.001 | 0.453 | 0.000 | 0.262 | 0.013 | 0.474 | 0.012 | 0.450 |
| VideoChat2 | - | 0.077 | 0.679 | 0.079 | 0.595 | 0.073 | 0.577 | 0.004 | 0.378 | 0.000 | 0.211 | 0.101 | 0.721 | 0.056 | 0.527 |
| GUI-Vid | - | 0.096 | 0.754 | 0.149 | 0.689 | 0.131 | 0.700 | 0.051 | 0.637 | 0.003 | 0.460 | 0.082 | 0.705 | 0.085 | 0.657 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | R. | 0.108 | 0.797 | 0.142 | 0.717 | 0.080 | 0.682 | 0.075 | 0.714 | 0.011 | 0.602 | 0.117 | 0.746 | 0.089 | 0.710 |
| E. | 0.099 | 0.768 | 0.136 | 0.700 | 0.075 | 0.655 | 0.066 | 0.695 | 0.011 | 0.592 | 0.113 | 0.743 | 0.083 | 0.692 | |
| Qwen-VL-Max | R. | 0.087 | 0.704 | 0.098 | 0.650 | 0.112 | 0.639 | 0.009 | 0.519 | 0.003 | 0.465 | 0.106 | 0.725 | 0.069 | 0.617 |
| E. | 0.075 | 0.638 | 0.095 | 0.647 | 0.094 | 0.600 | 0.009 | 0.512 | 0.009 | 0.475 | 0.103 | 0.712 | 0.064 | 0.597 | |
| H. | 0.080 | 0.632 | 0.083 | 0.589 | 0.092 | 0.617 | 0.013 | 0.520 | 0.007 | 0.452 | 0.099 | 0.703 | 0.062 | 0.585 | |
| GPT-4V | R. | 0.159 | 0.824 | 0.224 | 0.772 | 0.206 | 0.766 | 0.040 | 0.673 | 0.030 | 0.579 | 0.174 | 0.777 | 0.139 | 0.732 |
| E. | 0.149 | 0.813 | 0.201 | 0.752 | 0.207 | 0.746 | 0.035 | 0.659 | 0.017 | 0.566 | 0.160 | 0.762 | 0.128 | 0.716 | |
| H. | 0.156 | 0.805 | 0.205 | 0.745 | 0.203 | 0.748 | 0.034 | 0.644 | 0.025 | 0.559 | 0.159 | 0.763 | 0.130 | 0.711 | |
| GPT-4o | H. | 0.137 | 0.802 | 0.196 | 0.761 | 0.199 | 0.732 | 0.035 | 0.683 | 0.022 | 0.533 | 0.154 | 0.774 | 0.124 | 0.714 |
| ChatUnivi | - | 0.093 | 0.679 | 0.085 | 0.604 | 0.106 | 0.616 | 0.005 | 0.437 | 0.000 | 0.258 | 0.076 | 0.698 | 0.061 | 0.548 |
| Minigpt4Video | - | 0.026 | 0.547 | 0.026 | 0.513 | 0.035 | 0.548 | 0.001 | 0.411 | 0.000 | 0.236 | 0.015 | 0.529 | 0.017 | 0.464 |
| VideoChat2 | - | 0.089 | 0.683 | 0.078 | 0.605 | 0.061 | 0.555 | 0.002 | 0.355 | 0.000 | 0.190 | 0.086 | 0.710 | 0.053 | 0.516 |
| GUI-Vid | - | 0.114 | 0.725 | 0.144 | 0.693 | 0.123 | 0.700 | 0.048 | 0.641 | 0.002 | 0.518 | 0.083 | 0.686 | 0.085 | 0.661 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | E. | 0.089 | 0.771 | 0.189 | 0.704 | 0.189 | 0.710 | 0.023 | 0.619 | 0.016 | 0.570 | 0.149 | 0.749 | 0.109 | 0.687 |
| Qwen-VL-Max | R. | 0.041 | 0.640 | 0.084 | 0.528 | 0.066 | 0.549 | 0.008 | 0.484 | 0.004 | 0.445 | 0.089 | 0.673 | 0.049 | 0.553 |
| E. | 0.037 | 0.634 | 0.074 | 0.498 | 0.065 | 0.541 | 0.005 | 0.443 | 0.003 | 0.383 | 0.089 | 0.683 | 0.045 | 0.530 | |
| GPT-4V | R. | 0.106 | 0.809 | 0.242 | 0.757 | 0.210 | 0.733 | 0.029 | 0.653 | 0.028 | 0.619 | 0.170 | 0.763 | 0.131 | 0.723 |
| E. | 0.089 | 0.771 | 0.192 | 0.705 | 0.190 | 0.713 | 0.023 | 0.619 | 0.016 | 0.571 | 0.150 | 0.750 | 0.110 | 0.688 | |
| GPT-4o | H. | 0.075 | 0.809 | 0.241 | 0.755 | 0.188 | 0.719 | 0.038 | 0.677 | 0.014 | 0.581 | 0.137 | 0.747 | 0.116 | 0.715 |
| ChatUnivi | - | 0.076 | 0.675 | 0.079 | 0.588 | 0.096 | 0.594 | 0.007 | 0.482 | 0.001 | 0.368 | 0.063 | 0.670 | 0.054 | 0.563 |
| Minigpt4Video | - | 0.017 | 0.416 | 0.013 | 0.369 | 0.019 | 0.405 | 0.000 | 0.279 | 0.000 | 0.103 | 0.010 | 0.392 | 0.010 | 0.327 |
| VideoChat2 | - | 0.057 | 0.641 | 0.077 | 0.560 | 0.063 | 0.523 | 0.004 | 0.402 | 0.000 | 0.272 | 0.075 | 0.654 | 0.046 | 0.509 |
| GUI-Vid | - | 0.083 | 0.682 | 0.130 | 0.628 | 0.126 | 0.644 | 0.023 | 0.500 | 0.001 | 0.393 | 0.071 | 0.659 | 0.072 | 0.584 |
| Models | Setting | Static | Sequential | Prediction | Description | Caption | Conversation | Avg. | |||||||
| BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | BLEU | B.S. | ||
| Gemini-Pro-1.5 | R. | 0.113 | 0.739 | 0.126 | 0.693 | 0.086 | 0.658 | 0.061 | 0.685 | 0.012 | 0.586 | 0.090 | 0.674 | 0.081 | 0.673 |
| E. | 0.106 | 0.728 | 0.131 | 0.680 | 0.072 | 0.622 | 0.055 | 0.655 | 0.015 | 0.550 | 0.084 | 0.679 | 0.077 | 0.652 | |
| Qwen-VL-Max | R. | 0.079 | 0.599 | 0.076 | 0.591 | 0.080 | 0.595 | 0.002 | 0.444 | 0.006 | 0.370 | 0.072 | 0.666 | 0.053 | 0.544 |
| E. | 0.064 | 0.609 | 0.087 | 0.567 | 0.089 | 0.608 | 0.003 | 0.445 | 0.004 | 0.398 | 0.073 | 0.647 | 0.053 | 0.546 | |
| H. | 0.089 | 0.634 | 0.078 | 0.580 | 0.093 | 0.612 | 0.003 | 0.409 | 0.005 | 0.344 | 0.080 | 0.656 | 0.058 | 0.539 | |
| GPT-4V | R. | 0.172 | 0.800 | 0.186 | 0.737 | 0.212 | 0.745 | 0.040 | 0.671 | 0.021 | 0.592 | 0.145 | 0.728 | 0.129 | 0.712 |
| E. | 0.160 | 0.763 | 0.169 | 0.703 | 0.198 | 0.759 | 0.034 | 0.621 | 0.012 | 0.527 | 0.116 | 0.709 | 0.115 | 0.680 | |
| H. | 0.173 | 0.781 | 0.196 | 0.748 | 0.220 | 0.775 | 0.046 | 0.672 | 0.021 | 0.577 | 0.133 | 0.724 | 0.132 | 0.713 | |
| GPT-4o | H. | 0.156 | 0.792 | 0.185 | 0.754 | 0.213 | 0.769 | 0.040 | 0.683 | 0.019 | 0.588 | 0.121 | 0.717 | 0.122 | 0.717 |
| ChatUnivi | - | 0.076 | 0.628 | 0.063 | 0.573 | 0.103 | 0.605 | 0.009 | 0.413 | 0.000 | 0.191 | 0.057 | 0.643 | 0.051 | 0.509 |
| Minigpt4Video | - | 0.015 | 0.504 | 0.024 | 0.473 | 0.023 | 0.527 | 0.001 | 0.326 | 0.000 | 0.155 | 0.009 | 0.469 | 0.012 | 0.409 |
| VideoChat2 | - | 0.098 | 0.657 | 0.081 | 0.593 | 0.067 | 0.577 | 0.007 | 0.344 | 0.000 | 0.162 | 0.065 | 0.654 | 0.053 | 0.498 |
| GUI-Vid | - | 0.128 | 0.737 | 0.144 | 0.664 | 0.133 | 0.721 | 0.041 | 0.605 | 0.004 | 0.452 | 0.058 | 0.644 | 0.084 | 0.637 |
Appendix E Prompts
In this section, we provide detailed prompts for models and human annotators. Figure 19 shows the guideline of human annotation, Figure 18 shows the prompt for leveraging LLMs to refine grammarly mistakes and polish sentence for human annotations. Figure 20, Figure 21, and Figure 22 present the prompt for Human-MLLM collaboration method to generate GUI-orientaed tasks. Figure 23 illustrate the prompt for benchmarking MLLMs, different GUI scenarios and different QA type has different prompt. Figure 24 and Figure 25 show prompt for LLM-as-a-Judge for free-form as well as conversational tasks and multiple-choice QA respectively.
Appendix F Case Study
In this section, we provide detailed case studies for six GUI scenarios, each divided into two parts. Figure 26 and Figure 27 show example frames and various tasks associated with them. Figure 28 and Figure 29 for IOS, Figure 30 and Figure 31 for multiple-windows interaction, Figure 34 and Figure 35 for website, and Figure 36 and Figure 37 for XR respectively.