### v2

# GUI-World: A Video Benchmark and Dataset for Multimodal GUI-oriented Understanding

Abstract
Recently, Multimodal Large Language Models (MLLMs) have been used as agents to control keyboard and mouse inputs by directly perceiving the Graphical User Interface (GUI) and generating corresponding commands. However, current agents primarily demonstrate strong understanding capabilities in static environments and are mainly applied to relatively simple domains, such as Web or mobile interfaces. We argue that a robust GUI agent should be capable of perceiving temporal information on the GUI, including dynamic Web content and multi-step tasks. Additionally, it should possess a comprehensive understanding of various GUI scenarios, including desktop software and multi-window interactions. To this end, this paper introduces a new dataset, termed GUI-World, which features meticulously crafted Human-MLLM annotations, extensively covering six GUI scenarios and eight types of GUI-oriented questions in three formats. We evaluate the capabilities of current state-of-the-art MLLMs, including Image LLMs and Video LLMs, in understanding various types of GUI content, especially dynamic and sequential content. Our findings reveal that current models struggle with dynamic GUI content without manually annotated keyframes or operation history. On the other hand, Video LLMs fall short in all GUI-oriented tasks given the sparse GUI video dataset. Therefore, we take the initial step of leveraging a fine-tuned Video LLM, GUI-Vid, as a GUI-oriented assistant, demonstrating an improved understanding of various GUI tasks. However, due to the limitations in the performance of base LLMs, we conclude that using video LLMs as GUI agents remains a significant challenge. We believe our work provides valuable insights for future research in dynamic GUI content understanding. All the dataset and code are publicly available at: https://gui-world.github.io.
1 Introduction
Multimodal Large Language Models (MLLMs), such as GPT-4V(ision) (OpenAI, 2023) and LLaVA (Liu et al., 2023b), have significantly contributed to the development of both vision and language domains (Yin et al., 2024). These models bring forth innovative solutions and paradigms for traditional visual tasks, including visual reasoning (Yang et al., 2023b), medical image interpretation (Li et al., 2024b), and applications in embodied agents (Huang et al., 2023). One particularly promising area is Graphical User Interface (GUI) understanding, which holds significant potential for real-world applications, such as webpage comprehension (Hong et al., 2024, Lai et al., 2024) and navigation by GUI agents (Yang et al., 2023a, Niu et al., 2024, Wang et al., 2024). The key challenges of GUI understanding are twofold: effective GUI agents are expected to (1) possess a deep understanding of GUI elements, including webpage icons, text identified through Optical Character Recognition (OCR), and page layouts, and (2) exhibit an exceptional ability to follow instructions within GUI contexts, such as conducting searches through search engines.
Despite significant progress, as illustrated in Table 1, prior studies on GUI-related datasets and benchmarks suffer the following limitations: (1) Inability to Handle Dynamic Environments. Most studies primarily emphasize the static features of GUI scenarios, often overlooking the importance of enabling MLLMs to effectively handle dynamic information and sequential operations. For instance, an agent’s task performance can be disrupted by unexpected elements such as pop-up advertisements, underscoring a gap in handling dynamic sequential tasks. (2) Limited Scenarios. Current research is typically restricted to Web-based or mobile environments, which limits to assess their generalization ability and robustness across scenarios. For instance, GUI-oriented models may need to operate across diverse platforms such as Windows, macOS, Linux, iOS, Android, and eXtended Reality (XR) environments. Operations may sometimes involve multiple windows. Therefore, expanding research to encompass these varied environments will enhance the adaptability and effectiveness.
To mitigate these gaps, this paper introduces GUI-World, a comprehensive dataset containing 12,379 GUI videos, specifically designed to evaluate and enhance the capabilities of GUI agents. This dataset encompasses a wide range of GUI scenarios, including popular websites, desktop and mobile applications across various operating systems, multi-window interactions, as well as XR environments. The data collection process involves sourcing GUI videos from screen recordings and instructional videos on YouTube. Subsequently, we utilize a Human-MLLM collaborative approach to generate a diverse set of captions, complex queries, and multi-round conversation, ultimately constructing GUI-World.
Likewise, we also establish a comprehensive benchmark for GUI understanding, which encompasses nine mainstream MLLMs (e.g., GPT-4o (OpenAI, 2024) and Gemini-1.5-Pro (Team et al., 2023)), five keyframe selection strategies (e.g., UVD (Zhang et al., 2024d)), and six GUI scenarios, aiming to provide a thorough evaluation of the GUI-oriented understanding capabilities of MLLMs. As shown in Section 2, the assessment results indicate that most MLLMs struggle with GUI-World, highlighting their limited dynamic understanding of graphical interfaces and underscoring the need for further enhancement.
Using this dataset, we take the first step to fine-tune a GUI-oriented Video LLM that excels at handling dynamic and sequential GUI tasks, leading to substantial enhancements in the general capabilities and showcasing the utility and effectiveness of GUI-World. Additionally, we delve into discussing various factors critical to GUI understanding, including the integration of textual information, the number of keyframes, image resolutions, and vision perception, providing a pioneering and comprehensive study of the GUI domain.
Overall, the key contributions of this paper are threefold:
A New Dataset. We propose GUI-World, a comprehensive GUI dataset comprising 12,379 videos specifically designed to assess and improve GUI-oriented capabilities of MLLMs, spanning a range of categories and scenarios, including desktop, mobile, and XR environments. It stands as the first GUI-oriented instruction-tuning dataset in video domain.
Comprehensive Experiments and Valuable Insights. Our experiments indicate that most existing MLLMs continue to face challenges with GUI-oriented tasks, particularly in sequential and dynamic GUI content. Empirical results suggest that enhancing vision perception, such as more keyframes and higher resolution, can lead to substantial performance improvements in GUI tasks.
A Explorative GUI-oriented Video LLM. Based on GUI-World, we propose GUI-Vid, a GUI-oriented video LLM with enhanced capabilities to handle various and complex GUI tasks. GUI-Vid shows a significant improvement on the benchmark and achieves results comparable to the top-performing models, thereby paving ways for the future of GUI models.
| Dataset | Size | Sem. | VL | Video | Env Type | Task Coverage | Task | |||||
| Web. | Mob. | Desk. | XR | Seq. | Cro. | Dyn. | ||||||
| Rico (Deka et al., 2017) | 72,219 | Low | ✔ | ✔ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | UI Code/Layout Generation |
| MiniWoB++ (Liu et al., 2018) | 100 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | ✘ | Web Navigation |
| Screen2Words (Wang et al., 2021) | 22,417 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Summarization |
| MetaGUI (Sun et al., 2022) | 1,125 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Mobile Navigation |
| UGIF (Venkatesh et al., 2022) | 523 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Instruction Following |
| AITW (Rawles et al., 2023) | 715,142 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| Ferret-UI (You et al., 2024) | 123,702 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Grounding & Understanding |
| Spotlight (Li & Li, 2022) | 2.5M | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | GUI Understanding |
| WebArena (Zhou et al., 2023) | 812 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| Mind2Web (Deng et al., 2024) | 2,350 | Both | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| OmniAct (Kapoor et al., 2024) | 9,802 | Low | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | Code Generation |
| GUICourse (Chen et al., 2024c) | 10.7M | Both | ✔ | ✘ | ✔ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| MMINA (Zhang et al., 2024e) | 1,050 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✔ | ✘ | Web Navigation |
| AgentStudio (Zheng et al., 2024b) | 304 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| OSWorld (Xie et al., 2024) | 369 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| GUI-World (Ours) | 12,379 | Both | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | GUI Understanding |
| Instruction Following | ||||||||||||
2 GUI-World: A Dataset for GUI Understanding
2.1 Overview
We introduce GUI-World, a comprehensive dataset covering six GUI scenarios including video, human-annotated keyframes, as well as detailed captions and diverse types of QA produced by our data curation framework, aiming at benchmarking and enhancing the general GUI-oriented capabilities. These GUI scenarios encompass desktop operating systems (e.g., macOS, Windows) and mobile platforms (e.g., Android and iOS), websites, software, and even XR (e.g., GUI in Apple Vision Pro (Apple, 2024)). We divide the dataset into a train-test split, each containing 10,702 and 1,677 samples. Discussion for each scenario is referred to Appendix A.1.
As illustrated in Figure 3, the development of GUI-World follows a two-stage process. Further details about video and query statistics are outlined in Table 2, including distributions of keyframe counts, video durations, query lengths, and their corresponding golden answers. Figures 4 and 5 further illustrate these data distributions.
2.2 GUI Video Collection and Keyframe Annotation Process
We describe the pipeline for collecting screen recordings from student workers and GUI-related instructional videos from YouTube for GUI-World and the procedures followed to convert these videos into keyframe sequences.
| Category | Total Videos | Free-form | MCQA | Conversation | Total Frame. (Avg.) | Avg. Anno. |
| Software | 4,720 | 27,840 | 9,440 | 9,440 | 23,520 (4.983) | 7.558 |
| Website | 2,499 | 14,994 | 4,998 | 4,998 | 15,371 (6.151) | 6.862 |
| IOS | 492 | 2,952 | 984 | 984 | 2,194 (4.459) | 7.067 |
| Multi | 475 | 2,850 | 950 | 950 | 2,507 (5.277) | 7.197 |
| XR | 393 | 2,358 | 786 | 786 | 1,584 (4.030) | 10.970 |
| Android | 3,800 | 15,199 | 7,600 | 7,600 | 38,000 (10.000) | - |
| Summary | 12,379 | 76,673 | 24,758 | 24,758 | 83,176 (6.719) | 7.463 |
A significant portion of our video data is derived from screen recordings executed by student workers, which can directly reflect real-life GUI usage scenarios. A typical video collection scenario involves assigning a student worker a specific software task. The student begins by familiarizing themselves with the software, followed by recording a series of operations in a short video clip, such as “Sign up”, “Sign in”, “Create a New Page”, and “Invite Other Collaborators” in the software “Notion111https://www.notion.so/”.
Despite the high fidelity of these manually recorded videos, we encounter several challenges: (1) Student workers often require substantial time to acquaint themselves with professional software (e.g., MATLAB, Adobe After Effects (Ae)), which can hinder the progress of data collection. (2) The videos may lack comprehensiveness, typically capturing only commonly used operations and overlooking rarer functions crucial for dataset completeness. To address these issues, we also source videos from social media platforms that host a diverse array of GUI videos. Specifically, we download tutorial videos from YouTube—given its prevalence as a video-sharing platform—because they richly detail various GUI operations. These videos are then segmented into shorter clips, each representing a distinct sequence of operations.
The subsequent step involves annotating these video clips with keyframes and textual descriptions of each keyframe using custom-designed annotation software. Although several algorithms exist for keyframe extraction (Zhu et al., 2016, Yan et al., 2018, Mahasseni et al., 2017, OpenCV, ), they typically underperform with GUI videos where changes between frames might be minimal (e.g., a slight movement in the mouse cursor). To ensure high-quality datasets, we therefore perform manual extraction of these keyframes. Each keyframe is meticulously annotated to include details such as the operation performed, the purpose between two keyframes, the software or website used, mouse actions (e.g., scroll, click), and keyboard inputs (e.g., copy (Ctrl + C), paste (Ctrl + V), specific input). We detail our annotation process in Appendix A.3.
2.3 GUI Tasks Generation from Human-MLLM Collaboration
Drawing insights from prior research (Dekoninck et al., 2024), we develop a Human-MLLM collaboration pipeline to annotate captions and diverse types of QA specifically tailored for GUI comprehension. The process involves inputting an instructional prompt, a comprehensive description, key information (e.g., system or application), and a sequence of human-annotated keyframes into GPT-4V. As depicted in Table 11, GUI-World features various question types, detailed as follows:
Detailed and Summarized Captioning: This task challenges basic GUI knowledge and multimodal perception, also addressing the deficiency of detailed GUI content in video-caption pairs. Initially, GPT-4V generates two distinct descriptions for each video: one concentrating on fine-grained details and the other on overall information. Furthermore, GPT-4V provides a succinct summary, highlighting core operations and overarching objectives in the video.
Static GUI Content: This task challenges MLLM with textual, layout, and iconographic analysis of static GUI content. We instruct GPT-4V to generate free-form queries with a golden answer concerning static GUI elements or specific scenes that recur in more than two keyframes, ensuring their consistent presence in the video. Additionally, GPT-4V also crafts QA pairs that evaluate inferential skills in static content, focusing on interrelations among icons or textual information.
Dynamic and Sequential GUI Content: This task concentrates on temporal content in GUI video, such as dynamically changing interfaces, and aims to elucidate the sequential information and reasoning chains within GUI content. We direct GPT-4V to identify consistently changing elements to create queries for dynamic content. Moreover, predictive tasks are formulated on order and temporal relation in provided sequential images, challenging agents to anticipate future events or states.
In the last stage, human annotators will follow the guideline in Appendix A.3 and carefully review the entire video and MLLM-generated QA pairs to correct inaccuracies and hallucinations, as well as supplement information for both questions and answers to make these tasks more challenging.
3 Experiments and Analysis
3.1 Experimental Setups
Models.222 Given that GPT-4V is announced to be deprecated, we use GPT-4o for certain ablation studies to ensure that our results provide longer-term reference value. We conduct evaluations on five of the most popular vision LLMs: GPT-4V(ision) (OpenAI, 2023), GPT-4o (OpenAI, 2024), Qwen-VL-Max (Bai et al., 2023), LLaVA-OV-7B (Li et al., 2024a), and Gemini-Pro-1.5 (Team et al., 2023). Additionally, we test the effect of different vision inputs on GPT-4o, using no input, low and high-resolution settings, as well as without providing images, to further assess how resolution influences performance. Each model’s responses employ a three-step Chain-of-Thought (CoT) (Wei et al., 2022) process, i.e., “Describe-Analyze-Answer”, to evaluate their peak performance. Additionally, we assessed four advanced video LLMs—ChatUnivi (Jin et al., 2023), Minigpt4-video (Ataallah et al., 2024), Videochat2 (Li et al., 2023c), VideoLLaVA (Lin et al., 2023a) —for their performance on GUI content. See Appendix C for detailed setups.
Evaluation Metrics. To assess free-form questions and multiple-round conversations, we utilize the LLM-as-a-Judge methodology, which assigns a similarity score ranging from 1 to 5 between MLLM’s response and a predefined golden answer, already validated by previous studies (Zheng et al., 2023, Liu et al., 2023d, Chen et al., 2024a, Ye et al., 2024). For multiple-choice questions, we measure performance using accuracy as the primary evaluation metric.
Keyframe Extraction. We benchmark on three keyframe selection settings: (1) Linspave, where frames are evenly sampled at fixed time intervals within a video; (2) Program, with programmatic method Katna (KeplerLab, 2023); (3) Model-based, which leverages pre-trained vision representation from VIP (Ma et al., 2022) and R3M (Nair et al., 2022) to form UVD (Zhang et al., 2024d); and (4) Human, where humans select keyframes during the annotation process. We use Human setting for Image LLMs in our main experiment with an average of 6.719 frames (Table 2). For all other settings, we input 10 frames into each MLLM.
Additional Information Integration. To investigate the effectiveness of integrating image-caption models for LLMs—typically employed in natural videos—and the helpfulness of textual GUI content in accomplishing GUI-oriented tasks, we implement three experimental settings: Detailed Caption, Concise Caption, and Vision + Detailed Caption. GPT-4V is utilized to provide captions of these keyframes, integrating human annotators’ operational intents to more accurately describe each frame, being validated in Appendix A.3.
| Models | Software | Website | XR | Multi | IOS | Android | Avg. | ||||||||
| MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | ||
| Image LLMs | LLaVA-OV-7B | 56.9% | 2.641 | 48.4% | 2.588 | 59.6% | 2.709 | 52.9% | 2.306 | 58.1% | 2.717 | 24.3% | 2.675 | 50.0% | 2.606 |
| Gemini-Pro-1.5 | 82.9% | 3.385 | 79.2% | 3.412 | 83.3% | 3.108 | 83.4% | 3.246 | 80.3% | 3.467 | 78.5% | 3.168 | 81.3% | 3.298 | |
| Qwen-VL-Max | 75.8% | 2.651 | 75.5% | 2.698 | 77.6% | 2.373 | 66.9% | 2.490 | 74.3% | 2.633 | 74.2% | 2.559 | 74.0% | 2.568 | |
| GPT-4V | 86.0% | 3.520 | 79.8% | 3.655 | 83.4% | 3.265 | 76.9% | 3.449 | 79.9% | 3.453 | 81.3% | 3.466 | 81.2% | 3.469 | |
| GPT-4o | 86.5% | 3.644 | 83.3% | 3.740 | 84.3% | 3.285 | 81.1% | 3.654 | 83.3% | 3.558 | 90.0% | 3.561 | 84.8% | 3.573 | |
| Video LLMs | ChatUnivi | 28.4% | 2.389 | 22.2% | 2.349 | 20.6% | 2.161 | 17.5% | 2.275 | 22.6% | 2.337 | 23.0% | 2.390 | 22.4% | 2.317 |
| Minigpt4Video | 18.9% | 1.475 | 15.3% | 1.520 | 16.3% | 1.362 | 15.4% | 1.457 | 20.1% | 1.501 | 14.6% | 1.342 | 16.8% | 1.443 | |
| VideoChat2 | 45.5% | 2.144 | 42.6% | 2.221 | 44.0% | 2.005 | 40.4% | 2.222 | 40.2% | 2.169 | 44.7% | 2.119 | 42.9% | 2.147 | |
| Video-LLaVA | 52.9% | 2.290 | 52.4% | 2.410 | 44.2% | 2.258 | 45.9% | 2.329 | 49.7% | 2.319 | 51.3% | 2.259 | 49.4% | 2.311 | |
| GUI-Vid | 59.9% | 2.847 | 54.1% | 2.957 | 55.6% | 2.764 | 52.9% | 2.861 | 51.8% | 2.773 | 53.4% | 2.572 | 54.6% | 2.796 | |
| Models | Setting | Software | Website | XR | Multi | IOS | Android | Avg. | ||||||||
| Vision | Text | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |
| GPT-4V | ✘ | D.C. | 85.0% | 3.350 | 83.1% | 3.380 | 82.3% | 3.056 | 84.2% | 3.358 | 81.6% | 2.751 | 81.7% | 3.427 | 83.0% | 3.316 |
| ✘ | C.C. | 80.7% | 3.028 | 72.2% | 3.025 | 82.8% | 2.809 | 81.3% | 3.160 | 76.5% | 2.868 | 76.4% | 2.939 | 78.3% | 2.971 | |
| ✔ | D.C. | 82.5% | 3.494 | 83.2% | 3.682 | 85.9% | 3.191 | 83.9% | 3.617 | 80.9% | 3.516 | 84.9% | 3.758 | 83.5% | 3.543 | |
| Models | Caption | Complex Tasks | Conversation | Average | ||||
| Concise | Detailed | Static | Dyn. | Round 1 | Round 2 | |||
| Image LLMs | LLaVA-OV-7B | 2.149 | 1.762 | 1.868 | 2.448 | 2.947 | 3.492 | 2.641 |
| Gemini-Pro-1.5 | 3.306 | 3.035 | 2.945 | 3.093 | 3.573 | 3.790 | 3.298 | |
| Qwen-VL-Max | 2.474 | 1.711 | 2.137 | 2.433 | 3.223 | 3.257 | 2.651 | |
| GPT-4V | 3.352 | 2.509 | 3.053 | 3.229 | 3.928 | 4.163 | 3.520 | |
| GPT-4o | 4.048 | 3.028 | 3.125 | 3.340 | 4.129 | 4.318 | 3.644 | |
| Video LLMs | ChatUnivi | 1.587 | 1.240 | 1.705 | 2.090 | 2.698 | 3.366 | 2.389 |
| Minigpt4Video | 1.246 | 1.073 | 1.249 | 1.455 | 1.494 | 1.719 | 1.475 | |
| VideoChat2 | 1.992 | 1.312 | 1.812 | 1.920 | 2.342 | 2.720 | 2.144 | |
| Video-LLaVA | 1.519 | 1.241 | 1.657 | 1.959 | 2.587 | 3.293 | 2.290 | |
| GUI-Vid | 3.562 | 2.058 | 2.376 | 2.763 | 3.080 | 3.260 | 2.847 | |
3.2 Empirical Results
Commercial Vision LLMs outperform Open-source Video LLMs in Zero-shot Settings. Commercial vision LLMs, notably GPT-4V and GPT-4o, consistently outperform open-source video LLMs in zero-shot settings. As detailed in Table 3, GPT-4o exhibits superior performance across all GUI scenarios in complex tasks, reflected in its high scores in both multiple-choice and free-form queries, with an average of 84.8% and 3.573. Similarly, Gemini demonstrates strong capabilities in captioning and descriptive tasks within software and iOS environments, scoring 2.836 and 2.936, respectively, as shown in Table 26. Further analysis (Figure 6) reveals that GPT-4V excels in applications with minimal textual content and simple layouts, such as TikTok, health apps, and GitHub. In contrast, its performance drops in more intricate applications like Microsoft ToDo and XR software. As for video LLMs, their significantly poorer performance is attributed to two main factors: their inability to accurately interpret GUI content from user inputs and a lack of sufficient GUI-oriented pre-training, which is evident from their inadequate performance in basic captioning and description tasks. See Appendix D for other metrics and detailed fine-grained performance.
| Settings | Caption | Complex Tasks | Conversation | Average | |||
| Concise | Detailed | Static | Dyn. | Round 1 | Round 2 | ||
| Human | 3.911 | 3.031 | 3.131 | 3.318 | 3.981 | 4.132 | 3.573 |
| Program | 3.643 | 2.764 | 2.872 | 3.052 | 3.702 | 3.837 | 3.300 |
| Linspace | 3.749 | 2.941 | 3.000 | 3.077 | 3.687 | 3.843 | 3.440 |
| UVD+vip | 3.954 | 3.105 | 3.321 | 3.219 | 3.944 | 4.107 | 3.581 |
| UVD+r3m | 3.972 | 3.121 | 3.352 | 3.243 | 3.975 | 4.119 | 3.612 |
| Setting | Caption | Complex Tasks | Conversation | Average | |||
| Concise | Detailed | Static | Dyn. | Round 1 | Round 2 | ||
| w/o Vision | 2.187 | 1.872 | 2.486 | 2.979 | 3.760 | 4.059 | 2.891 |
| Low Resolution | 3.672 | 2.794 | 2.869 | 3.150 | 3.783 | 4.041 | 3.394 |
| High Resolution | 3.911 | 3.031 | 3.131 | 3.318 | 3.981 | 4.132 | 3.574 |
Dynamic GUI Tasks Continue to Challenge MLLMs. In the fine-grained tasks depicted in Table 5, GPT-4V and GPT-4o excel with dynamic GUI content and conversational tasks but struggle with providing detailed descriptions for entire videos and static content. This discrepancy is attributed to minor variations in GUI that significantly impact its semantic meaning. Enhancing the number of keyframes and the granularity of perception might mitigate these issues. Among video LLMs, ChatUnivi excels in conversational tasks by effectively leveraging contextual nuances, particularly in subsequent rounds, yet it underperforms in caption tasks. In contrast, GUI-Vid demonstrates proficiency in dynamic tasks but falls short in both captioning and static content. This gap is linked to deficiencies in backbone pretraining, which lacked comprehensive GUI content crucial for effective vision-text alignment, as evidenced by its poor performance in simple caption task shown in Table 26 and an instruction tuning process failed to fully address these shortcomings.
Vision Perception is Important to Dynamic GUI Content. As demonstrated in Table 7, integrating detailed textual information slightly outperforms purely vision-based inputs or detailed captions, akin to a Chain of Thought (CoT) (Wei et al., 2022) setting. Surprisingly, GPT-4V excels in caption tasks with just detailed captions, providing insights on enhancing specific GUI-oriented tasks through additional textual information. However, it still falls short in more challenging tasks, such as retrieving static or dynamic content. This underscores the critical role of visual perception in GUI environments, where even minor changes can significantly impact outcomes.
Keyframe Selection is Important for GUI-oriented Tasks. Our experiments demonstrate that model-based keyframe identifiers, originally developed for embodied AI applications, perform competitively with human-selected across both basic tasks (e.g., caption) and complex tasks (static and dynamic analysis). As shown in Table 6, GPT-4o exhibits significant performance improvements when utilizing these robotics-inspired model-based keyframe identifiers, with the UVD+VIP approach achieving optimal results. These findings suggest the potential to replace manual keyframe selection with automated approaches. Further analysis reveals that embodied AI keyframe identifiers successfully capture semantic transitions in GUI content, while linspace and program-based selection methods fail to do so, highlighting their particular suitability for GUI-oriented tasks. The substantial performance gaps observed between different selection methods underscore the critical importance of keyframe selection in this domain.
4 Exploring and Improving GUI-oriented Video LLMs
4.1 Method: Progressive Enhancement
We introduce our strategy to enhance the GUI-oriented capabilities of current MLLMs on both static and dynamic GUI content. Inspired by previous studies (Lai et al., 2024, Li et al., 2023b), we structure our methodology into two distinct fine-tuning stages, as illustrated in Figure 7. Initially, we fine-tune the MLLM on simpler tasks, such as description queries and captioning exercises, to instill a basic understanding of GUI elements. Subsequently, building on this foundation, the second stage aims to augment the MLLM’s proficiency with more complex and challenging tasks. Our fine-tuning is all based on the Supervised Fine-Tuning (SFT): , where is the input, is LLMs’ output, and denotes the model parameters that need to be optimized.
Stage-1: Learning Preliminary for GUI Content. The initial phase focuses on aligning GUI content with a pre-trained vision encoder and a base LLM, utilizing GUI videos accompanied by detailed descriptions and captions. This phase aims to embed a robust understanding of fundamental GUI concepts and terminology within the MLLM. By engaging the model in basically captioning various GUI components, the model learns to recognize and articulate the functionalities and visual characteristics of these elements, thereby laying a solid groundwork for GUI knowledge.
Stage-2: Mastering Advanced GUI Capability. Building on the foundational knowledge established in Stage 1, the second stage focuses on advancing the MLLM’s proficiency in interacting with GUI elements through more complex tasks. These tasks are designed to simulate real-world scenarios that the MLLM might encounter in GUI environments, which include predicting based on image sequences, engaging in conversations, retrieving both static and dynamic GUI elements, and performing reasoning tasks.
As illustrated in Figure 7, We employ the two-stage training architecture utilizing VideoChat2 (Li et al., 2023b) as our foundational model. Initially, videos and images are encoded using the UMT-L visual encoder (Li et al., 2023d). Subsequently, a QFormer compresses visual tokens into a smaller set of query tokens. Drawing inspiration from (Dai et al., 2023), we enhance the QFormer (Zhang et al., 2024c) by integrating instructions to enable it to extract visual representations pertinent to the given instructions. Additionally, we apply low-rank adaptation (LoRA (Hu et al., 2021)) to base LLM. This model is concurrently fine-tuned with the visual encoder and QFormer using a Vision-grounded Text Generation (VTG) loss: , where represents the visual tokens derived from the QFormer, and represents the text output grounded in the visual context.
4.2 Experiments
Experiment Setups. We use two dataset settings to fine-tune GUI-Vid, one with video only, and the other with video and image, detailed in Appendix C. We also vary the number of keyframes (8, 16) fed into GUI-Vid. All our experiments are conducted on A800 and 4090 GPUs.
Supreme Enhancement of GUI-Vid on Graphic-based Interface After Fine-tuning on GUI-World. As a pioneering study in training video LLMs as screen agents, GUI-Vid significantly outperforms the baseline model, showing an average improvement of 30% across various tasks and GUI scenarios, even surpassing the commercial vision LLM, Qwen-VL-Max. This enhancement is particularly notable in captioning and dynamic task that reason over image sequences, where GUI-Vid matches the performance of GPT-4V and Gemini-Pro. As depicted in Table 8, our two ablation studies during the fine-tuning phase demonstrate that utilizing GUI image-text captioning data significantly enhances the model’s preliminary understanding of GUI elements, outperforming training that relies solely on videos. Additionally, an increased number of keyframes correlates with improved performance across various scenarios, notably in environments featuring multiple windows and software applications. As shown in Figure 8, our two-stage progressive fine-tuning significantly enhances the performance in all GUI scenarios.
| Baseline | F.K. | E.K. | Data | Software | Website | XR | Multi | IOS | Android | Avg. | ||||||||
| I. | V. | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |||
| Baseline | - | 8 | - | - | 45.5% | 2.144 | 42.6% | 2.221 | 44.0% | 2.005 | 40.4% | 2.222 | 40.2% | 2.169 | 44.7% | 2.119 | 42.9% | 2.147 |
| - | 16 | - | - | 45.1% | 2.144 | 41.8% | 2.240 | 41.0% | 2.007 | 40.7% | 2.238 | 39.9% | 2.138 | 44.7% | 2.147 | 42.2% | 2.154 | |
| GUI-Vid | 8 | 8 | ✘ | ✔ | 58.3% | 2.709 | 53.6% | 2.817 | 62.2% | 2.626 | 54.2% | 2.627 | 53.1% | 2.708 | 54.9% | 2.501 | 56.0% | 2.665 |
| ✔ | ✔ | 59.9% | 2.856 | 54.1% | 2.925 | 59.0% | 2.751 | 52.1% | 2.837 | 50.0% | 2.756 | 54.0% | 2.571 | 54.8% | 2.782 | |||
| 16 | ✘ | ✔ | 59.0% | 2.709 | 55.1% | 2.821 | 62.8% | 2.645 | 53.3% | 2.624 | 55.5% | 2.727 | 55.7% | 2.501 | 56.9% | 2.671 | ||
| ✔ | ✔ | 59.9% | 2.847 | 54.1% | 2.957 | 55.6% | 2.764 | 52.9% | 2.861 | 51.8% | 2.772 | 53.4% | 2.572 | 54.6% | 2.796 | |||
Correlation between GUI Understanding and Other Mainstream GUI Tasks. In our explorative experiments, GUI-Vid still fails in some GUI operating tasks via code generation, which is due to the baseline LLM’s weak performance and the challenges of code generation instruction fine-tuning. To further demonstrate how GUI understanding capability enhances mainstream GUI-related tasks, including generating operational code (Cheng et al., 2024) and providing chat assistance (Hong et al., 2024), we conduct experiments as follows (detailed in Appendix D):
We conduct human study across 180 videos across 6 scenarios, where annotators will choose preferred responses from two response of different models when acting as GUI assistant (Table 17).
5 Related Work
MLLM-based Agents for GUI. Building upon the significant advancements in LLMs (Achiam et al., 2023, Meta, 2023a; b, ai, 2024, Huang et al., 2025) and advanced modality-mixing technologies (Li et al., 2023a, Alayrac et al., 2022), groundbreaking MLLMs such as GPT-4V (OpenAI, 2023) and Gemini-Pro (Team et al., 2023), along with open-source MLLMs like the LLaVA-1.6 series (Liu et al., 2023b; a), CogVLM (Wang et al., 2023b), and Qwen-VL series (Bai et al., 2023), have shown outstanding performance across various tasks (Yu et al., , Liu et al., 2023e, Chen et al., 2024b, Wu et al., 2023, Wake et al., 2023, Zhang et al., 2024b, Zhao et al., 2024, Gui et al., 2024). Venturing beyond text and single image, several studies are now exploring the integration of video modalities for tasks requiring dynamic or sequential visual content (Jin et al., 2023, Li et al., 2023b, Maaz et al., 2023, Lin et al., 2023a). In the GUI domain, leveraging the robust vision perception capabilities of MLLMs, applications such as WebAgents (Hong et al., 2024, Zhang et al., 2024a, Zheng et al., 2024b) and Mobile Agents (Wang et al., 2023a, You et al., 2024, Wang et al., 2021, Li et al., 2020b) have gained popularity for handling everyday tasks like navigation and VQA. Frontier research is also investigating the use of MLLMs as general control agents, such as in playing computer games (Tan et al., 2024, Lin et al., 2023b) and serving as OS copilots (Song et al., 2024, Xie et al., 2024), paving the way for more complex GUI operations.
GUI Benchmark & Dataset. Building upon the foundational work of Rico (Deka et al., 2017), the first mobile GUI video dataset, and AitW (Rawles et al., 2023), which features 715k episodes of sequential images, research has extensively covered mobile (Sun et al., 2022, Li et al., 2020a, Zhang et al., 2023) and web GUI environments (Lù et al., 2024, Zhou et al., 2023, Yao et al., 2022, Koh et al., 2024, Liu et al., 2024b). Mind2Web (Deng et al., 2024) stands out in web-based datasets with over 2,000 tasks from 137 websites across 31 domains. Advances continue into desktop GUIs with new toolkits (Zheng et al., 2024b), benchmarks (Kapoor et al., 2024, Mialon et al., 2023), and frameworks (Zheng et al., 2024a, Liu et al., 2023c, Niu et al., 2024). Research on GUI also transfers from comprehending single images in a static workspace (Hong et al., 2024) to sequential operations or multi-hop scenarios (Xie et al., 2024, Zhang et al., 2024e), challenging the understanding and operation capability of these powerful models.
6 Conclusion
In this paper, we have introduced GUI-World, a comprehensive GUI-oriented video dataset designed to benchmark and enhance understanding of virtual interfaces, especially sequential and dynamic tasks. This dataset extensively covers six scenarios and various tasks, addressing the previous research gap in comprehensively evaluating models’ capabilities in graphic-based understanding. We conduct extensive benchmarks on leading MLLMs and the first video-LLM-based assistant GUI-Vid fine-tuned on GUI-World specifically for GUI-oriented content, achieving results comparable to top-performing models, providing detailed insights into enhancing GUI-related capabilities. We believe our work offers valuable insights for future research in dynamic GUI content understanding.
Acknowledgements. Pan Zhou is partially supported by National Natural Science Foundation of China (NSFC) under grant No. 62476107. Dongping Chen and Yao Wan are supported by the Fundamental Research Funds for the Central Universities (HUST: 62400001). We would like to thank Yinuo Liu, Zhengyan Fu, Shilin Zhang, Yu, Tianhe Gu, Haokuan Yuan, and Junqi Wang for their insightful feedback and support.
Appendix A Details of Dataset Construction
A.1 Six Main GUI Categories
In earlier endeavors pertaining to GUI, such as those involving GUI testing (Kousar et al., 2023, Jorge et al., 2014, Kulesovs, 2015), the focus was segmented into GUIs for Website, Software, IOS and Android platforms. However, as a comprehensive GUI dataset, we include all potential GUI scenarios in our dataset to ensure that our data is the most comprehensive knowledge that the GUI agent needs to learn; we divide these scenarios into six categories:
Android. This category focuses on the GUI scenarios that occur within the Android operating system, which is predominantly used on smartphones. Android’s ubiquity in the mobile market has led to a wide variety of GUI designs and interaction patterns, making it a rich field for study. This category has been the subject of extensive scrutiny in scholarly works such as (Deka et al., 2017, Li et al., 2020a, Rawles et al., 2023, Cheng et al., 2024).
Software. This category encapsulates the GUI scenarios arising within software applications, whether they are standalone programs or components of a larger suite. The diversity of software applications, from productivity tools to creative suites, offers a wide range of GUI scenarios for exploration. The literature is rich with research in this area, such as (Zhan et al., 2024).
Website. This category is concerned with the GUI scenarios that manifest within a web browser. Given the ubiquity of web browsing in modern digital life, this category holds significant relevance. It holds a substantial representation in academic literature, with pioneering papers such as (Deng et al., 2024, Kapoor et al., 2024) proposing excellent GUI datasets for websites.
IOS. This category zeroes in on the GUI scenarios that transpire within the iOS operating system, the proprietary system for Apple devices like the iPhone and iPad. The iOS platform is known for its distinct design aesthetics and interaction patterns, providing a unique context for GUI research. A number of studies, such as (Beltramelli, 2018, Yan et al., 2023) make use of GUI information in IOS.
Multi-Windows. This category is dedicated to GUI scenarios that necessitate simultaneous interaction with multiple windows, a common occurrence in desktop environments where users often juggle between several applications or documents. Despite the common use of multi-window interaction in everyday GUI usage, there has been relatively little research into this area (Nakajima et al., 2013). The need for efficient multitasking in such scenarios presents unique challenges and opportunities for GUI design and interaction research. As of our knowledge, there are no specific datasets catering to these multi-window GUI scenarios.
XR. XR encompasses Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR) (Rauschnabel et al., 2022). Given the advancements in XR technology and the growing accessibility of commercial-grade head-mounted displays (Apple, 2024, Met, ), XR has emerged as a novel medium for human-computer interaction. This necessitates the exploration of GUI within XR environments. In these scenarios, the GUI takes on a 3D, immersive form (Sanders et al., 2019), demanding the agent to comprehend and navigate a 3D space. The emerging field of XR presents a new frontier for GUI research, with unique challenges and opportunities due to its immersive and interactive nature. To date, as far as we are aware, there are no datasets that specifically address GUI in the realm of XR.
A.2 Selected Website/Software
In our study, we select a diverse range of websites and software to comprehensively evaluate GUI understanding capabilities across various user scenarios. These selections cover essential categories such as social media, productivity tools, online shopping, and educational platforms, providing a broad spectrum of GUI environments.
Figure 10 shows an overall distribution and our selected app, software, and website in GUI-World.
These selections ensure that our study encompasses a wide array of user interactions and GUI complexities, thereby providing a robust evaluation of the current state-of-the-art methods in GUI understanding by MLLMs and comprehensively constructing a high-quality dataset.
A.3 Human Keyframes Annotation Process
Annotator’s Information The annotation is conducted by 16 authors of this paper and 8 volunteers independently. As acknowledged, the diversity of annotators plays a crucial role in reducing bias and enhancing the reliability of the benchmark. These annotators have knowledge in the GUI domain, with different genders, ages, and educational backgrounds. The education backgrounds of annotators are above undergraduate. To ensure the annotators can proficiently mark the data, we provide them with detailed tutorials, teaching them how to use software to record videos or edit video clips. We also provide them with detailed criteria and task requirements in each annotation process.
Recording Video. For self-recording videos, we employ OBS333https://obsproject.com/ on the Windows system for screen capturing and the official screen recording toolkit on the Mac/IOS system. This process necessitates human annotators to execute a series of targeted actions within specific websites or applications, which are subsequently captured as raw video footage. We provide a list of software and website for each annotator to first get familiar with then record video that operating on them. For some popular software or website such as chrome, we ask several annotator to record video of it. These actions, commonplace in everyday usage, enhance the reliability of our dataset. Subsequently, the raw videos are segmented into sub-videos, each encapsulating multiple actions (e.g., clicking a button) to achieve a specific objective (e.g., image search). The videos are then processed to extract keyframes annotated with detailed descriptions.
Edition Based on YouTube Videos. For sourcing videos from YouTube, we utilize a search protocol formatted as “[website name/application name] + tutorial” to compile relevant video lists. Human annotators first review these videos to understand the primary operations they depict. These videos are then divided into sub-videos, each containing several actions directed towards a single goal (e.g., image search). Like the self-recorded footage, these segments are processed to isolate keyframes and furnish them with descriptive annotations.
Keyframes Annotation. After obtaining the GUI video clips, human annotators will filter out the keyframes of the operations based on the video content and the mouse and keyboard actions at that time. They will also label the sub-operations or targets between the two keyframes. Once the annotation is complete, the annotators will provide an overall description of the entire video, summarizing the main goal of the human operations in the video. After all the information is annotated, we will use an LLM to refine the text content, reducing any errors made by human annotators and adjusting the sentence structure. The prompt we use for the LLM to polish the human annotations is shown in Figure 11 and Figure 12.
Human-LLM Cooperated Instruction Generation. To curate and refine the golden answer of each video-instruction pair generated by GPT-4V, given that the raw response from GPT-4V may contain harmful content or hallucinations. The role of humans in the golden answer generation process is to enhance the difficulty of the questions and remove harmful and incorrect content, as shown in Table 10.
| Raw response from GPT-4V | Golden answer from humans |
| “Q”: “What subcategories are visible under the ‘Lifestyle’ section in the static GUI content of the video?” | “Q”: “What subcategories are visible under the ‘Lifestyle’ section in the static GUI content of the video?” |
| “A”: “In the ‘Lifestyle’ section of the telegram.com website, the visible subcategories are ‘Worcester Living’, ‘Faith’, and ‘Food’. These are clearly displayed in a horizontal list just below the main ‘Lifestyle’ category header, indicating different areas of content that users can explore within the lifestyle section.” | “A”: “The ‘Lifestyle’ section of the telegram.com website features visible subcategories such as ‘Worcester Living’, ‘Faith’, and ‘Food’. These are prominently shown in a horizontal list beneath the main ‘Lifestyle’ header, offering users various content areas to explore within the section.” |
Human Verifying GPT-4V Annotated Captions. We evaluate the quality of annotations from GPT-4V by selecting 1,000 detailed descriptions and captions generated by GPT-4V, which are then assessed by human annotators. The high satisfaction rate of 98% underscores the quality and relevance of the GPT-4V annotations.
| T. | Question | Examples |
| Caption | Detailed | Q: Please provide a detailed description of what occurs throughout these sequential GUI images. |
| Description | A: The video shows a user taking the 16 Personalities test on a Windows desktop using the Edge browser… | |
| Summarized | Q: Write a clear description of the video, make sure the key features are well covered. | |
| Caption | A: Creating a new IT team in Todoist by selecting industry, job function, role, team size, and inviting members. | |
| Static | Layout, | Q: What related searches are suggested on the right side of the Bing results for ’emnlp 2024’? |
| Icon Retrieval | A: The suggested related searches shown include ’emnlp 2024 miami’, ’eacl 2024 call for papers’… | |
| Textual | Q: What is the estimated time to complete the content for Week 2 of the course? | |
| Retrieval | A: The estimated time to complete the content for Week 2 of the course is 1 hour… | |
| Interrelations | Q: What is the name of the browser and the tab where the user performs the product search? | |
| in GUI Content | A: The browser is Microsoft Edge, and the user performs the product search in the eBay tab. | |
| Dynamic | Content | Q: What specific action does the user take after turning their head to the left to view the left side of the page? |
| Retrieval | A: After turning their head to the left to view the left side of the page, the user performs… | |
| Prediction | Q: Given the mouse is over ’Add NeurIPS 2024 DB Track Submission,’ what’s the likely next step? | |
| A: It would be to click on the ’Add NeurIPS 2024 Datasets and Benchmarks Track Submission’ button… | ||
| Sequential | Q: Scrolls down from the ‘Moon Gravity’, which of the following cheats? A. Change Weather B. Skyfall … | |
| Reasoning | A: [[B]] |
Appendix B Dataset Analysis
In this section, we provide an analysis of the length distribution of QA in each GUI scenario, as illustrated in Figure 13 and Figure 14. Questions focused on sequential and predictional tasks are slightly longer than other types, while the golden answer of static tasks tends to be longer. Length of Question-answer pair in various GUI scenarios is similarly distributed, with questions in Android environment being slightly shorter, and answers in XR environment being longer.
Appendix C Details of Experiments Setups
C.1 Fine-tune dataset construction
We use two settings to fine-tune GUI-Vid, one with video-text pairs only, and the other with video-text and image-text pairs, which are all GUI content:
Video Only. In this setting, we only train GUI-Vid with video-text pairs in GUI-World, as shown in Table 12.
Video-Image. Inspired by the pre-trained process of Videochat2, we include image-text pairs to help the visual encoder align GUI knowledge. These images are selected from our GUI-World, MetaGUI (Sun et al., 2022), and OmniAct (Kapoor et al., 2024) for high-quality GUI content. Subsequently, we use GPT-4V to generate a detailed description and a concise caption for each image. Finally, we construct a dataset consisting of video-text and image-text pairs for gaining comprehensive GUI-oriented capabilities.
| Stage | Data types | Amount |
| 1 | Detailed Description | 14,276 |
| Concise Caption | 7,138 | |
| 2 | GUI VQA | 21,414 |
| Multiple-Choice QA | 14,276 | |
| Conversation | 7,138 |
| Stage | Data types | Source | Type | Amount |
| 1 | GUI-World | Video | Detailed Description | 14,276 |
| Concise Caption | 7,138 | |||
| Image | Detailed Description | 5,555 | ||
| Concise Caption | 5,555 | |||
| MetaGUI | Image | Detailed Description | 19,626 | |
| Concise Caption | 19,626 | |||
| OmniAct | Detailed Description | 260 | ||
| Concise Caption | 260 | |||
| 2 | GUI-World | Video | GUI VQA | 21,414 |
| Multiple-Choice QA | 14,276 | |||
| Conversation | 7,138 |
C.2 Hyperparameter Settings
In this section, we will introduce the hyperparameters of MLLMs to facilitate experiment reproducibility and transparency. We divide them into three parts: the inference phase during benchmark and dataset construction, the LLM-as-a-Judge phase, and the fine-tuning phase. All our experiments were conducted on a server equipped with dual A800 and dual 4090 GPUs.
Inference. We empirically study 7 MLLMs, involving 4 Image LLMs and 3 Video LLMs, with their hyperparameters detailed as follows:
Gemini-Pro-1.5 (Team et al., 2023): We use the default settings, which set temperature as 0.4, top-p as 1, and max-token as 2048. It should be noted that during our project, Gemini-Pro-1.5 is still under the user request limit, which only provides 100 requests per day, making our benchmark difficult. Given that Gemini hasn’t launched Pay-as-you-go444https://ai.google.dev/pricing, we will include benchmark results on ‘Human’ setting as soon as possible.
Qwen-VL-Max (Bai et al., 2023): We use the default settings for Qwen-VL-Max, with top-p as 0.8 and max-token as 2048. Given that the input context window is merely 6,000 for Qwen, we scale the resolution for all images to 0.3.
ChatUnivi (Jin et al., 2023): We use ChatUnivi-7B built upon Vicuna-v0-7B and set the max frame as 100, temperature as 0.2, and max-token as 1024.
Minigpt4video (Ataallah et al., 2024): We use the suggested settings555https://github.com/Vision-CAIR/MiniGPT4-video for this model and the max-frame are set as 45, with only the max-token being modified to 1024.
VideoChat2 & GUI-Vid (Li et al., 2023c): For a fair comparison, we set the same hyperparameters for VideoChat2 & GUI-Vid. We set the max-token as 1024, top- as 0.9, temperature as 1.0, max-frame as 8/16, repetition penalty as 1.2, and length penalty as 1.2.
LLM-as-a-Judge. We investigate four LLM-as-a-Judge in giving a similarity score for the MLLM’s response and ground truth, namely GPT-4 (Achiam et al., 2023), ChatGPT (OpenAI, 2023), LLaMA-3-70b-instruct (Meta, 2023b), and Mixtral-8x22b-instruct-v0.1 (ai, 2024). Hyperparameter settings are detailed as follows:
GPT-4 & ChatGPT. We set the temperature as 0.6 and others as default.
LLaMA-3-70b-instruct. We set the temperature as 0.6, top-p as 0.9, top-k as 50.
Mixtral-8x22b-instruct-v0.1. We set top-p as 0.7, top-k as 50, and temperature as 0.7.
Fine-tune. We include several hyperparameter settings in experiment settings and ablation studies, as shown in Table 14.
| Config | Setting |
| input frame | 8 |
| input resolution | 224 |
| max text length | 512 |
| input modal | I. + V. |
| optimizer | AdamW |
| optimizer momentum | |
| weight decay | 0.02 |
| learning rate schedule | cosine decay |
| learning rate | 2e-5 |
| batch size | 4 |
| warmup epochs | 0.6 |
| total epochs | 3 |
| backbone drop path | 0 |
| QFormer drop path | 0.1 |
| QFormer dropout | 0.1 |
| QFormer token | 96 |
| flip augmentation | yes |
| augmentation | MultiScaleCrop [0.5, 1] |
| Models | Pearson() | Spearman() | Kendall() | $ per Benchmark() |
| GPT-4 | 0.856 | 0.853 | 0.793 | $ |
| ChatGPT | 0.706 | 0.714 | 0.627 | 12$ |
| Llama-3-70b-instruct | 0.774 | 0.772 | 0.684 | 12$ |
| Mixtral-8x22b-instruct-v0.1 | 0.759 | 0.760 | 0.670 | $ |
C.3 Evaluation.
Given the complexity of free-form answers in GUI scenarios, the evaluation includes specific positions of GUI elements, textual content, and comparing the response to the golden answer. LLM-as-a-Judge has been widely used in previous studies for complex evaluation tasks (Zheng et al., 2023, Liu et al., 2023d). Therefore, we leverage LLM-as-a-Judge (Zheng et al., 2023) in a similar setting to MM-vet (Yu et al., ), which compares the MLLM’s response to the golden answer. We carefully evaluate the accessibility of leveraging LLM-as-a-Judge, selecting 1,000 samples covering 6 free-form questions mentioned in our dataset. As shown in Table 15, GPT-4 outperforms other LLMs, exhibiting a better human alignment on providing a similarity score for the response compared to the golden answer, although it is approximately 10 times more expensive than other models.
Appendix D Additional Experiments Results
Removing Data from MetaGUI and OmniAct. As shown in Table 16, after excluding OmniAct and MetaGUI from the training data, the model showed performance decreases in certain areas while demonstrating notable improvements in others. Specifically, we observed significant enhancements in MCQA tasks and XR-related capabilities. Conversely, models trained with MetaGUI and OmniAct exhibited superior performance on general free-form questions. These experimental results validate that models trained exclusively on GUI-World can achieve highly competitive performance, thereby confirming the exceptional quality of our dataset.
| Models | Software | Website | XR | Multi | IOS | Android | Avg. | |||||||||
| MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |||
| VideoChat2 | 45.50% | 2.144 | 42.60% | 2.221 | 44.00% | 2.005 | 40.40% | 2.222 | 40.20% | 2.169 | 44.70% | 2.119 | 42.90% | 2.147 | ||
| GUI-Vid | 59.90% | 2.847 | 54.10% | 2.957 | 55.60% | 2.764 | 52.90% | 2.861 | 51.80% | 2.773 | 53.40% | 2.572 | 54.60% | 2.796 | ||
59.70% | 2.811 | 57.30% | 2.844 | 59.62% | 2.853 | 50.00% | 2.799 | 55.31% | 2.769 | 57.67% | 2.547 | 56.60% | 2.767 | ||
| Scenarios | GUI-Vid | Tie | VideoChat2 |
| Software | 82.7% | 13.3% | 4.0% |
| Website | 86.0% | 12.0% | 2.0% |
| XR | 88.0% | 8.7% | 3.3% |
| Multi | 85.3% | 10.0% | 8.7% |
| IOS | 92.0% | 6.0% | 2.0% |
| Android | 82.0% | 16.0% | 2.0% |
| Average | 86.0% | 11.0% | 3.7% |
| Model | GUI-World | VisualAgentBench | VideoGUI | OS-World |
| GPT-4o | 1 | 1 | 1 | 2 |
| GPT-4V | 2 | 2 | 2 | 1 |
| Gemini-1.5-Pro | 3 | 3 | 3 | 3 |
| Qwen-VL-Max | 4 | 4 | 4 | / |
In this section, we first provide an ablation study on keyframe selection methods. Then, we conduct statistics and human preference experiments on correlations of GUI understanding capability to other mainstream GUI-related tasks. Furthermore, we provide detailed, performance on newly released models after our submission of the first version, followed by very detailed results on each task in each GUI scenario.
Ablation Study on Keyframe Identify Methods. Firstly, we show performance on model-based keyframe identify methods in Table 6, with details of UVD+VIP and UVD+R3M in Tables 24 and 25.
Correlation between GUI Understanding and Other Mainstream GUI Tasks. Furthermore, we conduct additional analysis and experiments to show how GUI understanding capability helps mainstream GUI-related tasks, including generating code to operate GUI (Cheng et al., 2024) and assist people through chat (Hong et al., 2024). Both demonstrate the strong correlation between GUI understanding capability and specific tasks for GUI agents.
We compare the benchmark results on GUI-World with existing benchmarks (Xie et al., 2024, Qinghong Lin et al., 2024, Liu et al., 2024c) for operating on GUI as shown in Table 18, and find that the results generally match, i.e., the stronger the understanding ability, the stronger the agent performance.
For the definition of chat helping humans, we select 180 videos from the benchmark, choosing 30 videos for each scenario. We ask 5 human annotators to pose the question they most wanted to ask after watching each video. We then use GUI-Vid, both before and after fine-tuning, to answer these questions. The human annotators who ask the questions are then asked to indicate which answer is more helpful. The results are shown in Table 17, demonstrating that models trained in GUI understanding are more favored by people when acting as GUI agents.
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| XR | 0.442 | 1.100 | 2.686 | 2.055 | 1.808 | 1.654 | 2.258 |
| Android | 0.513 | 1.162 | 2.952 | 1.858 | 1.673 | 1.763 | 2.259 |
| IOS | 0.497 | 1.143 | 2.966 | 1.992 | 1.680 | 1.654 | 2.319 |
| Multi | 0.459 | 1.106 | 2.863 | 2.069 | 1.781 | 1.772 | 2.329 |
| Website | 0.524 | 1.183 | 3.059 | 2.102 | 1.736 | 1.371 | 2.410 |
| Software | 0.529 | 1.241 | 2.942 | 1.958 | 1.657 | 1.519 | 2.290 |
| Average | 0.494 | 1.156 | 2.911 | 2.005 | 1.722 | 1.622 | 2.311 |
| Scenarios | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| XR | 0.596 | 1.867 | 3.123 | 2.580 | 2.147 | 1.987 | 2.709 |
| Android | 0.243 | 1.675 | 3.338 | 2.360 | 1.980 | 2.189 | 2.675 |
| IOS | 0.581 | 1.762 | 3.229 | 2.536 | 2.051 | 2.017 | 2.717 |
| Multi | 0.355 | 1.069 | 2.982 | 2.437 | 1.870 | 2.541 | 2.541 |
| Website | 0.484 | 1.729 | 3.123 | 2.422 | 1.854 | 2.004 | 2.588 |
| Software | 0.569 | 1.762 | 3.220 | 2.448 | 1.868 | 2.149 | 2.641 |
| Average | 0.471 | 1.644 | 3.169 | 2.464 | 1.961 | 2.148 | 2.645 |
Performance of Newly Released Models in GUI-World Test Set. We evaluate two latest models, LLaVA-Next-Video-7B-DPO (Liu et al., 2024a) and Video-LLaVA (Lin et al., 2023a). We show their performance in Tables 20 and 19. Our model outperforms these in most tasks, except conversation, likely due to their use of DPO during training.
Performance on MV-Bench. We conducted evaluations on MVBench (Li et al., 2023c) across four models, as shown in Table 21. While our model showed some performance degradation compared to the original model, it still significantly outperformed the other two baseline models. We would like to note that the primary purpose of training GUI-Vid was to demonstrate the effectiveness of our dataset, rather than achieving state-of-the-art performance. Therefore, we did not employ the strongest baseline model or optimal training methodologies, which may have contributed to some catastrophic forgetting. However, given that GUI-Vid is specifically designed for GUI-related tasks, we believe this performance trade-off is acceptable within our research context.
| Category | VideoChat2 | GUI-Vid | VideoLLaMA | VideoChatGPT |
| Action Prediction | 47.5 | 39.0 | 25.5 | 26.0 |
| Unexpected Action | 60.0 | 57.0 | 39.0 | 26.5 |
| Object Existence | 58.0 | 54.0 | 48.0 | 54.0 |
| Object Interaction | 71.5 | 51.0 | 40.5 | 28.0 |
| Object Shuffle | 41.0 | 30.5 | 38.0 | 40.0 |
| Moving Direction | 23.0 | 21.0 | 22.5 | 23.0 |
| Action Localization | 23.0 | 30.5 | 22.5 | 20.0 |
| Scene Transition | 88.0 | 65.5 | 43.0 | 31.0 |
| Action Count | 39.5 | 34.5 | 34.0 | 30.5 |
| Moving Count | 42.0 | 29.0 | 22.5 | 25.5 |
| Moving Attribute | 58.5 | 53.5 | 32.5 | 39.5 |
| State Change | 44.5 | 41.0 | 45.5 | 48.5 |
| Character Order | 36.5 | 39.0 | 32.5 | 29.0 |
| Egocentric Navigation | 35.0 | 34.0 | 40.0 | 33.0 |
| Episodic Reasoning | 38.5 | 37.0 | 30.0 | 29.5 |
| Average | 47.1 | 41.1 | 34.4 | 32.3 |
Performance on Mind2Web-Multimodal. We evaluated our model on the Mind2Web-Multimodal dataset (Deng et al., 2024) to demonstrate the effectiveness of fine-tuning on GUI-World. Mind2Web-Multimodal is a multiple-choice GUI benchmark that assesses GUI operation capabilities, where each sample comprises an image, a task description, and response options. Our evaluation included four experimental settings: zero-shot inference using both VideoChat2 and GUI-Vid, as well as versions of these models fine-tuned on Mind2Web-Multimodal. We conducted the fine-tuning process for 3 epochs. As shown in Tables 22 and 23, fine-tuning on GUI-World enhances the models’ GUI understanding capabilities and improves their performance on operational benchmarks.
| Category | ZeroShot | Fine-tuned | ||
| VideoChat2 | GUI-Vid | VideoChat2 | GUI-Vid | |
| Cooking | 21.28% | 23.40% | 23.40% | 26.24% |
| Education | 17.61% | 16.61% | 24.92% | 28.24% |
| Finance | 26.02% | 29.59% | 26.02% | 32.65% |
| Government | 18.80% | 18.28% | 26.89% | 27.42% |
| Health | 17.08% | 17.37% | 22.58% | 26.48% |
| Home service | 19.08% | 20.14% | 21.55% | 28.27% |
| Housing | 12.21% | 14.13% | 18.20% | 21.84% |
| Job | 16.25% | 18.00% | 20.00% | 20.25% |
| Moving | 14.36% | 17.44% | 20.00% | 24.62% |
| Pet | 14.05% | 17.97% | 21.57% | 21.24% |
| Shipping | 13.42% | 12.99% | 20.78% | 20.34% |
| Social media | 22.15% | 24.37% | 22.78% | 29.43% |
| Weather | 27.14% | 22.14% | 26.43% | 20.00% |
| Overall | 17.53% | 18.59% | 22.37% | 25.14% |
| Category | ZeroShot | Fine-tuned | ||
| VideoChat2 | GUI-Vid | VideoChat2 | GUI-Vid | |
| Auto | 20.00% | 16.00% | 19.00% | 18.00% |
| Department | 9.52% | 11.90% | 14.29% | 16.67% |
| Digital | 19.73% | 19.73% | 15.65% | 20.41% |
| Event | 34.65% | 25.74% | 21.78% | 22.77% |
| General | 18.32% | 18.85% | 17.80% | 20.94% |
| Music | 16.67% | 15.66% | 21.69% | 24.10% |
| Other | 17.81% | 21.92% | 27.40% | 35.62% |
| Restaurant | 10.70% | 14.44% | 17.65% | 16.58% |
| Sports | 16.98% | 18.87% | 20.75% | 26.42% |
| Overall | 17.94% | 17.96% | 18.84% | 21.20% |
Additional Experiment Results on GUI-World. For captioning tasks, Table 26 shows comprehensive experimental results among six scenarios. For scores of LLM-as-a-Judge in a specific task, see Tables 27, 28, 29, 30, and 31. For performance in fine-grain (application level), see Figure 16 for Gemini-Pro and Figure 17 for Qwen-VL-Max.
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| Software | 86.2% | 3.297 | 4.282 | 3.354 | 3.478 | 4.112 | 3.749 |
| Website | 82.0% | 3.248 | 4.155 | 3.415 | 3.567 | 4.074 | 3.744 |
| XR | 84.2% | 2.980 | 3.775 | 3.034 | 3.122 | 3.587 | 3.347 |
| Multi | 82.1% | 3.391 | 4.165 | 3.466 | 3.404 | 3.868 | 3.659 |
| IOS | 86.0% | 3.157 | 4.017 | 3.353 | 3.492 | 4.050 | 3.648 |
| Mobile | 80.7% | 2.827 | 3.871 | 2.970 | 3.014 | 3.844 | 3.340 |
| Average | 83.5% | 3.150 | 4.044 | 3.265 | 3.346 | 3.923 | 3.581 |
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| Software | 85.8% | 3.290 | 4.273 | 3.352 | 3.458 | 4.134 | 3.741 |
| Website | 82.7% | 3.282 | 4.114 | 3.460 | 3.591 | 4.065 | 3.746 |
| XR | 87.7% | 3.010 | 3.861 | 3.142 | 3.161 | 3.600 | 3.433 |
| Multi | 83.6% | 3.237 | 4.129 | 3.503 | 3.417 | 3.897 | 3.737 |
| IOS | 86.4% | 3.165 | 4.094 | 3.328 | 3.480 | 4.078 | 3.663 |
| Android | 80.6% | 2.835 | 3.876 | 2.968 | 3.072 | 3.865 | 3.353 |
| Average | 84.5% | 3.136 | 4.058 | 3.292 | 3.363 | 3.940 | 3.612 |
| Models | Setting | Software | Website | XR | Multi | IOS | Android | Avg. | |||||||
| Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | Cap. | Des. | ||
| Gemini-Pro-1.5 | R. | 3.659 | 2.837 | 3.613 | 2.860 | 2.995 | 2.590 | 3.276 | 2.470 | 3.678 | 2.936 | - | - | 3.444 | 2.739 |
| E. | 3.350 | 2.468 | 3.159 | 2.422 | 2.837 | 2.279 | 2.824 | 2.109 | 3.394 | 2.519 | 3.185 | 2.312 | 3.125 | 2.351 | |
| Qwen-VL-Max | R. | 2.381 | 1.758 | 2.326 | 1.681 | 2.172 | 1.772 | 2.035 | 1.463 | 2.513 | 1.662 | 2.141 | 1.565 | 2.261 | 1.650 |
| E. | 2.459 | 1.693 | 2.317 | 1.599 | 2.167 | 1.638 | 2.190 | 1.438 | 2.189 | 1.615 | 2.002 | 1.429 | 2.221 | 1.569 | |
| H. | 2.474 | 1.711 | 2.457 | 1.698 | 2.383 | 1.777 | 1.910 | 1.346 | 2.577 | 1.795 | 2.474 | 1.711 | 2.360 | 1.665 | |
| GPT-4V | R. | 3.579 | 2.676 | 3.612 | 2.699 | 2.975 | 2.525 | 3.281 | 2.661 | 3.757 | 2.775 | 3.655 | 2.755 | 3.479 | 2.682 |
| E. | 3.141 | 2.301 | 3.293 | 2.380 | 2.471 | 2.085 | 3.063 | 2.324 | 3.624 | 2.611 | 3.201 | 2.312 | 3.132 | 2.335 | |
| H. | 3.352 | 2.509 | 3.702 | 2.750 | 3.050 | 3.556 | 3.524 | 2.673 | 3.670 | 2.588 | - | - | 3.460 | 2.614 | |
| GPT-4o | H. | 4.048 | 3.028 | 4.067 | 3.233 | 3.398 | 2.729 | 3.869 | 3.111 | 4.014 | 2.993 | 4.071 | 3.095 | 3.911 | 3.869 |
| ChatUnivi | - | 1.587 | 1.240 | 1.569 | 1.254 | 1.417 | 1.148 | 1.575 | 1.267 | 1.480 | 1.146 | 1.778 | 1.249 | 1.568 | 1.217 |
| Minigpt4Video | - | 1.246 | 1.073 | 1.200 | 1.057 | 1.320 | 1.106 | 1.130 | 1.034 | 1.190 | 1.076 | 1.184 | 1.061 | 1.212 | 1.068 |
| VideoChat2 | - | 1.992 | 1.312 | 1.817 | 1.307 | 1.838 | 1.426 | 2.222 | 1.433 | 2.169 | 1.270 | 2.119 | 1.294 | 1.900 | 1.340 |
| GUI-Vid | - | 3.562 | 2.085 | 3.655 | 2.167 | 3.747 | 2.153 | 3.370 | 1.742 | 3.566 | 2.071 | 2.662 | 1.248 | 3.427 | 1.911 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 3.279 | 3.050 | 3.560 | 3.579 | 3.796 | 3.452 |
| E. | 2.983 | 2.491 | 3.432 | 3.405 | 3.760 | 3.215 | |
| Qwen-VL-Max | R. | 2.317 | 2.271 | 2.802 | 2.995 | 3.069 | 2.656 |
| E. | 2.256 | 2.198 | 2.821 | 2.861 | 3.144 | 2.627 | |
| H. | 2.308 | 2.078 | 2.832 | 3.061 | 3.358 | 2.698 | |
| GPT-4V | R. | 3.461 | 3.214 | 3.754 | 3.778 | 4.029 | 3.648 |
| E. | 3.197 | 2.808 | 3.487 | 3.717 | 3.954 | 3.433 | |
| H. | 3.498 | 3.255 | 3.727 | 3.731 | 4.061 | 3.655 | |
| C.C. | 1.746 | 2.738 | 3.645 | 3.363 | 3.632 | 3.025 | |
| D.C. | 2.704 | 2.917 | 3.686 | 3.680 | 3.901 | 3.380 | |
| H.+D.C. | 3.313 | 3.221 | 3.852 | 3.850 | 4.171 | 3.682 | |
| GPT-4o | H. | 3.443 | 3.373 | 3.672 | 4.086 | 4.122 | 3.740 |
| ChatUnivi | - | 1.701 | 1.668 | 2.524 | 2.514 | 3.338 | 2.349 |
| Minigpt4Video | - | 1.309 | 1.233 | 1.766 | 1.439 | 1.854 | 1.520 |
| VideoChat2 | - | 1.771 | 1.777 | 2.288 | 2.461 | 2.812 | 2.221 |
| GUI-Vid | - | 2.406 | 2.341 | 3.544 | 3.135 | 3.355 | 2.957 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 2.892 | 2.505 | 3.543 | 3.222 | 3.611 | 3.154 |
| E. | 2.814 | 2.163 | 3.510 | 3.108 | 3.455 | 3.006 | |
| Qwen-VL-Max | R. | 2.047 | 1.968 | 2.712 | 2.879 | 3.132 | 2.469 |
| E. | 2.125 | 1.973 | 2.658 | 2.760 | 3.029 | 2.499 | |
| H. | 1.886 | 1.920 | 2.656 | 2.727 | 3.012 | 2.373 | |
| GPT-4V | R. | 2.934 | 2.668 | 3.392 | 3.291 | 3.714 | 3.200 |
| E. | 2.222 | 2.153 | 3.310 | 3.151 | 3.618 | 2.892 | |
| H. | 2.893 | 2.778 | 3.538 | 3.364 | 3.747 | 3.265 | |
| C.C. | 1.744 | 2.412 | 3.327 | 3.080 | 3.485 | 2.809 | |
| D.C. | 2.427 | 2.409 | 3.518 | 3.176 | 3.749 | 3.056 | |
| H.+D.C. | 2.775 | 2.635 | 3.580 | 3.235 | 3.734 | 3.191 | |
| GPT-4o | H. | 2.871 | 2.745 | 3.370 | 3.596 | 3.836 | 3.285 |
| ChatUnivi | - | 1.660 | 1.420 | 2.205 | 2.250 | 3.270 | 2.161 |
| Minigpt4Video | - | 1.225 | 1.161 | 1.610 | 1.347 | 1.465 | 1.362 |
| VideoChat2 | - | 1.654 | 1.547 | 2.192 | 2.099 | 2.529 | 2.005 |
| GUI-Vid | - | 2.444 | 2.147 | 3.347 | 2.836 | 3.036 | 2.764 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 2.538 | 2.410 | 3.296 | 3.152 | 3.402 | 2.959 |
| E. | 2.545 | 2.049 | 2.972 | 2.930 | 3.389 | 2.777 | |
| Qwen-VL-Max | R. | 1.793 | 1.872 | 2.770 | 2.897 | 3.122 | 2.432 |
| E. | 1.866 | 1.780 | 2.730 | 2.627 | 3.105 | 2.362 | |
| H. | 1.884 | 1.969 | 2.913 | 2.689 | 3.104 | 2.490 | |
| GPT-4V | R. | 3.185 | 2.655 | 3.745 | 3.699 | 3.973 | 3.452 |
| E. | 2.902 | 2.406 | 3.636 | 3.420 | 3.729 | 3.219 | |
| H. | 3.000 | 2.952 | 3.801 | 3.597 | 3.889 | 3.449 | |
| C.C. | 2.097 | 2.973 | 3.774 | 3.331 | 3.621 | 3.160 | |
| D.C. | 2.671 | 2.979 | 3.849 | 3.466 | 3.822 | 3.358 | |
| H.+D.C. | 3.037 | 3.162 | 4.079 | 3.748 | 4.036 | 3.617 | |
| GPT-4o | H. | 3.108 | 3.106 | 3.829 | 4.043 | 4.188 | 3.654 |
| ChatUnivi | - | 1.658 | 1.623 | 2.514 | 2.384 | 3.199 | 2.275 |
| Minigpt4Video | - | 1.205 | 1.186 | 1.690 | 1.400 | 1.801 | 1.457 |
| VideoChat2 | - | 1.754 | 1.774 | 2.479 | 2.420 | 2.699 | 2.222 |
| GUI-Vid | - | 2.485 | 2.067 | 3.537 | 2.954 | 3.247 | 2.861 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | R. | 3.076 | 2.637 | 3.370 | 3.366 | 3.615 | 3.213 |
| E. | 2.852 | 2.356 | 3.137 | 3.126 | 3.566 | 3.007 | |
| Qwen-VL-Max | R. | 2.438 | 2.244 | 2.923 | 3.102 | 3.273 | 2.779 |
| E. | 2.303 | 2.150 | 2.614 | 3.145 | 3.264 | 2.659 | |
| H. | 1.884 | 1.969 | 2.913 | 2.689 | 3.104 | 2.490 | |
| GPT-4V | R. | 3.364 | 3.080 | 3.684 | 3.766 | 4.184 | 3.614 |
| E. | 3.209 | 2.774 | 3.545 | 3.611 | 4.006 | 3.427 | |
| H. | 3.107 | 2.830 | 3.631 | 3.680 | 4.011 | 3.453 | |
| C.C. | 1.788 | 2.291 | 3.511 | 3.212 | 3.542 | 2.868 | |
| D.C. | 2.751 | 2.732 | 3.654 | 3.642 | 3.842 | 3.324 | |
| H.+D.C. | 3.090 | 2.965 | 3.740 | 3.786 | 3.994 | 3.516 | |
| GPT-4o | H. | 3.183 | 2.993 | 3.460 | 4.050 | 4.141 | 3.558 |
| ChatUnivi | - | 1.771 | 1.642 | 2.408 | 2.559 | 3.307 | 2.337 |
| Minigpt4Video | - | 1.291 | 1.219 | 1.698 | 1.556 | 1.737 | 1.501 |
| VideoChat2 | - | 1.955 | 1.803 | 2.145 | 2.315 | 2.626 | 2.169 |
| GUI-Vid | - | 2.262 | 2.133 | 3.401 | 2.843 | 3.224 | 2.773 |
| Models | Setting | Static | Sequential | Prediction | Conversation1 | Conversation2 | Average |
| Gemini-Pro-1.5 | E. | 2.703 | 2.460 | 3.157 | 3.642 | 3.881 | 3.168 |
| Qwen-VL-Max | R. | 1.887 | 1.804 | 2.398 | 2.823 | 3.056 | 2.309 |
| E. | 1.785 | 1.630 | 2.311 | 2.605 | 3.233 | 2.277 | |
| GPT-4V | R. | 3.116 | 3.047 | 3.477 | 3.924 | 4.008 | 3.515 |
| E. | 2.705 | 2.470 | 3.175 | 3.647 | 3.885 | 3.176 | |
| C.C. | 2.092 | 2.243 | 3.139 | 3.443 | 3.782 | 2.939 | |
| D.C. | 3.015 | 2.890 | 3.357 | 3.883 | 3.990 | 3.427 | |
| GPT-4o | H. | 3.057 | 3.220 | 3.373 | 3.981 | 4.186 | 3.561 |
| ChatUnivi | - | 1.835 | 1.654 | 2.317 | 2.712 | 3.433 | 2.390 |
| Minigpt4Video | - | 1.183 | 1.159 | 1.507 | 1.342 | 1.521 | 1.342 |
| VideoChat2 | - | 1.732 | 1.754 | 2.125 | 2.340 | 2.645 | 2.119 |
| GUI-Vid | - | 2.010 | 1.928 | 3.053 | 2.755 | 3.105 | 2.572 |
Appendix E Prompts
In this section, we provide detailed prompts for models and human annotators. Figure 19 shows the guideline of human annotation, and Figure 18 shows the prompt for leveraging LLMs to refine grammarly mistakes and polish sentence for human annotations. Figures 20, 21, and 22 present the prompt for Human-MLLM collaboration method to generate GUI-orientaed tasks. Figure 23 illustrate the prompt for benchmarking MLLMs, different GUI scenarios and different QA type has different prompt. Figure 24 presents the prompt for free-form QA using LLM-as-a-Judge. Figure 25 presents the prompt for multiple-choice QA.
Appendix F Case Study
In this section, we provide detailed case studies for six GUI scenarios, each divided into two parts. The following examples show human-annotated frames and various tasks associated with them: